### v1

# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities). We then decompose visual reasoning process into two stages: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features multi-run proactive perception and decoupled vision-reasoning capabilities. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on a wide range of benchmarks for both open-source and closed-source models. In addition, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
1 Introduction
In recent years, large language models (LLMs) (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to integrate visual understanding with textual reasoning simultaneously due to inherent modality differences and deficient training data. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2 illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
To address the challenges, a promising solution is to extract visual information from images into textual form to assist LVLMs in reasoning (Yin et al., 2023b; Mitra et al., 2024). For example, Ghosh et al. (2024) instruct LVLMs to generate fine-grained captions to facilitate the subsequent reasoning process. However, these existing methods exhibit two primary limitations: (i) The visual extraction process is question-agnostic and reasoning-free, that the image description is not targeted for a given question, and reasoning is not involved to deduce extra information for better descriptions. This drawback, termed as “passive”, results in irrelevant or insufficient information, and subsequent performance degradation. (ii) The reasoning process entangles visual understanding with textual reasoning abilities of a single LVLM. As mentioned above, the limited capabilities of LVLMs do not always successfully handle both abilities and produce high-quality reasoning. Hence, both of these limitations hinder the performance of LVLMs in multi-modal reasoning tasks.
To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features proactive (i.e., question-oriented and reasoning-involved) visual extraction and disentangled vision-reasoning capabilities. Specifically, we first decouple multi-modal reasoning capacity into two sub-tasks: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom). The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions. As illustrated in Figure 1, our framework can be divided into three stages: Action, Judgment, and Summary, with each involving different agents. In the Action stage, a Dispatcher selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive more insights. Unlike passive methods, all three agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency. Subsequently, a Referee in the Judgment stage evaluates the sufficiency of known information, deciding whether to return to the Action stage or proceed to the Summary stage, where all information is consolidated for a Summarizer to generate the final answer. Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby addressing the limitations of LVLMs.
Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models. Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2%. Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent, as well as the advantages of proactive information acquisition. With the verified superiority over the simultaneous inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
The main contributions of this work are threefold:
  1. We propose a novel multi-modal reasoning framework named ProReason, featuring iterative proactive perception and decoupled vision-reasoning capabilities.
  2. Empirically, we identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
  3. ProReason showcases the remarkable feasibility of integrating LLMs for boosted visual reasoning, illuminating the potential for LLM-assisted LVLM reasoning in future research.
2 Preliminary Observations
We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities. All experiments are conducted with three recent LVLMs for robustness: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023). Further details are provided in Sec. 4.1.
2.1 Challenges in Visual Reasoning: Limitations of LVLMs
Chain-of-Thought(CoT) (Wei et al., 2022) has been extensively verified to enhance the performance of LLMs. Here, we explore its impact on the reasoning performance of LVLMs. Similarly, we require the models to “think step by step” before generating final answers. As shown in Table 1, a counter-intuitive phenomenon is observed: compared to “Direct” answering method, the introduction of CoT consistently incurs slight performance degradation across all three models. Inspired by the findings of Ghosh et al. (2024) that LVLMs often rely on prior language knowledge yet neglecting visual information, we hypothesize that the degraded performance might be caused by the image-irrelevant reasoning of CoT.
To validate the hypothesis, we design a pipeline referencing Liu et al. (2023c) to measure the correlation between model responses and image contents. Specifically, given a question, we take the reasoning process generated by GPT-4o-mini [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence, whose superiority on MMMU task is verified in Sec. 4.2, as the golden solution. Subsequently, a more powerful judge (i.e., GPT-4 (Achiam et al., 2023)) is employed to evaluate the relevance of responses of candidate models (e.g., Llama3-LLaVA-NeXT-8B) and this golden solution with the prompt template in Figure 5, and gives a comprehensive score from 1 to 5, with 5 indicating the highest relevance. Since the golden solution contains the essential information needed to solve the problem within the image, a higher correlation indicates a stronger relevance to the image content. As shown in Table 2, the reasoning chains of incorrect responses are significantly less relevant to the images than those of correct answers. An exemplary case is shown in Figure 2, where Llama3-LLaVA-NeXT-8B attempts to use the right-hand rule to solve the task, but its reasoning process is almost unrelated to the image, ultimately leading to an incorrect conclusion. This highlights the limitation of CoT for multi-modal reasoning tasks, that image-irrelevant reasoning process provokes performance degradation, since the limited capabilities of LVLMs cannot effectively handle both image and text information simultaneously.
Table 1: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
Method Model
Llama3-LLaVA-NeXT-8B LLaVA-OneVision-Qwen2-7B-OV Qwen-VL-Chat
Direct 41.8 49.1 36.2
CoT 41.5 46.5 35.4
VDGD 42.3 49.7 37.1
Table 2: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
Model Image Relevance Score
True False
Llama3-LLaVA-NeXT-8B 3.70 3.36
LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
Qwen-VL-Chat 3.63 2.80
Table 3: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion. “True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
Score Llama3-LLaVA-NeXT-8B
True False
Detail Level 4.43 3.93
Question Relevance 3.87 3.30
Reasoning Effective Info Inclusion 3.91 3.57
2.2 Drawbacks of Passive Information Extraction
Passive visual reasoning techniques suffer insufficient and irrelevant visual information, despite mitigating the oversight of images by converting them into detailed captions, like Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024). To support this claim, we generate fine-grained image captions using GPT-4o-mini with the prompt shown in Figure 6. We then incorporate these captions into the prompts for LVLMs to facilitate the reasoning process. As listed in Table 1, while these image descriptions improve the performance of LVLMs, the gains are marginal, consistently amounting to less than 1%. This underscores the limited utility of captions generated by passive methods.
For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information necessary for reasoning, respectively. Similar to Sec. 2.1, we pair the image captions with the reasoning process of GPT-4o-mini, and employ GPT-4 as a judge to rate the captions from 1 to 5 across these dimensions. The prompt template is shown in Figure 5. As shown in Table 3, the captions for correct responses of Llama3-LLaVA-NeXT-8B receive higher scores across all three criteria, highlighting the importance of better captions for multi-modal reasoning. Additionally, all captions score significantly lower in the Question Relevance and Reasoning Effective Info Inclusion dimensions than the Detail Level dimension, indicating that while the captions are detailed, they often lack relevance to the questions. Figure 2 shows a case where Llama3-LLaVA-NeXT-8B utilizes fine-grained captions to solve a question from the MMMU benchmark. As illustrated, although the caption exhaustively describes the image content, it incorrectly describes the wires in the image as octagons, and misses information about the locations of these wires. This information is irrelevant to the target question, thus offering minimal assistance to LVLMs. In summary, our analysis highlights the drawbacks of passive visual reasoning enhancement techniques in terms of information insufficiency and redundancy, due to their question-agnostic property.
(1) ProReason
(2) Passive Visual Reasoning Enhancement Technique
(3) Multi-step Reasoning Framework for LLMs
Figure 1: Overview and comparison of workflows of ProReason, VDGD and ReAct. Unlike existing works (e.g., VDGD and ReAct), our proposed method enables the model to actively acquire necessary information from the images, making it more suitable for LVLMs.
3 Method
As illustrated in Figure 1, ProReason consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component. The entire workflow comprises three steps: Action, Judgment, and Summary, effectively integrating the proactive visual understanding (i.e., eyesight) and textual reasoning (i.e., wisdom) capacities for enhanced multi-modal reasoning performance.
3.1 Action Step
Action step is the core of question-oriented visual information extraction, driven by three sub-agents: Dispatcher, Vision Expert, and Reasoning Expert. The Dispatcher orchestrates the workflow, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more. The responses from both the Vision Expert and the Reasoning Expert are stored in a textual Memory component.
Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Reasoning Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty). The Dispatcher then generates a query $q$ for the chosen expert. If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{v}$, which is then stored in the Memory. When the Reasoning Expert is selected, it provides a response consisting of the reasoning process and the final answer $A_{r}$ based on the query $q$ and known information in the Memory, before only $A_{r}$ is stored in Memory. Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT and ReAct, thereby suffering less from redundant information.
3.2 Judgment Step
Judgment step is conducted by a sub-agent called Referee, which evaluates whether the information stored in Memory is sufficient to answer the question $Q$. The input to the Referee includes the question $Q$ and available information in the Memory. If the Memory contains adequate information to answer the question $Q$, the Referee outputs the identifier “SOLVABLE”; otherwise, it outputs “UNSOLVABLE”. If the Referee’s output is SOLVABLE, the workflow precedes to the Summary phase. Conversely, if the output is UNSOLVABLE, the Action phase is re-executed to gather more necessary information. This allows the Referee to collaborate closely with the three sub-agents in the Action step, enabling the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods.
3.3 Summary Step
The Summary step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$. This step is mainly powered by a sub-agent called Summarizer. Once the Referee determines that the information in the Memory is sufficient to address the question $Q$, and outputs the identifier SOLVABLE, the Summarizer will be called to draw a conclusion, based on the Memory information. This conclusion represents the answer to the question $Q$ of the whole framework, and will be evaluated by the performance metrics.
3.4 Advantages of ProReason
Reduced information Mission or Redundancy.
With the close collaboration of the Dispatcher, Vision Expert, Reasoning Expert, and Referee agents, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy. Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions. This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
LLM-assisted Multi-modal Reasoning.
In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designate pipeline to complete the tasks. This decomposition-before-integration approach not only circumvents the inherent differences between modalities, but also allows visually irrelevant sub-agents to be performed by text-only LLMs. Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power multi-modal reasoning process for performance enhancement.
4 Experiments
In this section, we first evaluate the performance of our ProReason framework against recent baselines on multiple benchmarks, followed by an in-depth ablation analysis of different components.
4.1 General Setup
Datasets.
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly as follows:
  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities. Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. [^2] To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. For the results using the official MME calculation method, please refer to Table 9, Table 10, Table 11, and Table 12.
  • MMMU evaluates multi-modal models with multidisciplinary tasks that require college-level domain-specific knowledge and detailed reasoning. It comprises 11,500 questions across 30 disciplines and 183 sub-fields, emphasizing advanced perception and domain-specific reasoning.
  • MathVista focuses on more challenging mathematical reasoning tasks that demand precise visual recognition and compositional reasoning. It includes 6,141 examples from 31 multi-modal mathematics datasets.
  • HallusionBench evaluates models’ ability to reason with images such as statistical charts, emphasizing nuanced visual understanding. It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
Base Models.
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources. As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API. In contrast, Llama3-LLaVA-NeXT-8B is a fully open-sourced LVLM developed by the LLaVA team based on the Llama-3-8B LLM (Dubey et al., 2024) and CLIP vision encoder (Radford et al., 2021). Qwen2.5-72B-Instruct is a robust LLM designed to deliver high-quality instruction following and handle complex tasks. It will be utilized for achieving LLM-assisted multi-modal reasoning.
Baselines.
Besides the most basic method where models are instructed to answer questions directly, we compare ProReason with two categories of peer methods. To determine the benefits of proactive information extraction in ProReason, we first consider two SOTA passive visual reasoning methods, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024). Additionally, we choose two multi-step reasoning frameworks of LLMs, Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022), to demonstrate the effects of directly migrating LLM solutions to LVLMs.
  1. Direct. As indicated by the name, models are required to answer questions directly without dedicated prompts. This baseline is set to evaluate the initial performance of base models.
  2. VDGD. With image caption prefixed to text instruction, VDGD selects tokens that deviate the least from the description using a formula based on Kullback-Leibler divergence.
  3. CCoT. Given an image and the question, CCOT first generates a scene graph of the image with LVLMs, and then extracts the answer by prompting the LVLMs with the graph.
  4. CoT. CoT is an advanced prompting method that encourages LLMs to break complex tasks down into a series of easy steps, which has been applied broadly and verified to boost the reasoning performance remarkably (Chu et al., 2023).
  5. ReAct. ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors. To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Reasoning Experts, respectively. This aligns with our notions for easy understanding, and is shown in Figure 1.
Implementation Details.
The prompt templates for all methods are shown in Figures 6, 7 and 8. Specifically for ProReason, in order to prevent an infinite loop, if Dispatcher selects Vision Expert or Reasoning Expert to obtain information up to 5 times, and Referee still judges that the existing knowledge in Memory is not enough to solve the question, the information collection phase will be terminated immediately, and then Summarizer will directly answer the question based on the available information in Memory. This setup, refined through multiple trials, is the most effective. Additionally, since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini.
Table 4: Performance of multiple approaches with two base models across four visual reasoning benchmarks. “Hallu.” is the abbreviation of HallusionBench. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
Model Method Dataset
MME MMMU MathVista Hallu.
Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
ProReason 71.5 (+10.0) 52.5 (+10.7) 38.8 (+1.7) 50.9 (+5.1)
Average 66.30 44.22 36.06 44.16
GPT-4o-mini Direct 79.2 48.4 53.0 56.0
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
ProReason 90.4 (+11.2) 61.6 (+13.2) 54.9 (+1.9) 58.9 (+2.9)
Average 84.63 54.82 52.63 55.23
Table 5: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks. The red and blue texts indicate the improvements and reductions relative to the case of using the same method without the assistance of text-only LLMs, respectively.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
ReAct 68.5 46.7 31.7 43.6
ProReason 71.5 52.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 71.0 (+2.5) 50.4 (+3.7) 34.6 (+2.9) 40.4 (-3.2)
ProReason 81.3 (+9.8) 56.8 (+4.3) 48.8 (+10.0) 52.3 (+1.4)
GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
ProReason 84.7 (+13.2) 54.5 (+2.0) 41.7 (+2.9) 53.1 (+2.2)
4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 4, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently outperforms all the other baselines for both Llama3-LLaVA-NeXT-8B and GPT-4o-mini model across all benchmarks, with a peak improvement as 13.2%, demonstrating the superiority and task robustness of ProReason.
Proactive information acquisition surpasses SOTA passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Sec. 2.2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. As listed in Table 4, when both capabilities are utilized concurrently, CoT consistently degrades performance compared to the “Direct” method across all benchmarks with Llama3-LLaVA-NeXT-8B model, consistent with the findings in Sec. 2.1. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, demonstrating the effectiveness of capability decoupling.
ProReason outperforms ReAct with even less token consumption. Given the similarity in multi-step reasoning, we compare ProReason and ReAct in terms of performance and cost. Specifically, as an LLM-specific multi-step reasoning framework, ReAct only outperforms the “Direct” method on MME and MMMU, but underperforms on MathVista and HallusionBench, showing inferior performance compared to the consistent improvements of ProReason. Furthermore, we compare their average token consumption of GPT-4o-mini. As shown in Table 8 of Appendix A, ProReason consumes significantly fewer tokens than ReAct on both MME and MathVista tasks, indicating its higher token efficiency and the importance of a compact Memory in reducing token usage. Coupled with better performance, this suggests the superiority of ProReason over LLM-specific ReAct framework.
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Sec. 3.4, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert as Llama3-LLaVA-NeXT-8B, and replace other agents with text-only LLMs. As listed in Table 5, with the assistance of either Qwen2.5-72B-Instruct or GPT-4o-mini, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
4.3 Relative Importance of Sub-agents
To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Reasoning Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini. Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is. Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage. The experimental results are presented in Table 6.
Table 6: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Dispatcher Vision Expert Reasoning Expert Referee Summarizer
MME 90.4 88.8 (-1.6) 84.7 (-5.7) 88.7 (-1.7) 89.6 (-1.1) 84.2 (-6.2)
MMMU 61.6 60.9 (-0.7) 54.5 (-7.1) 60.2 (-1.4) 51.5 (-10.1) 51.0 (-10.6)
Summarizer is the most crucial sub-agent, closely followed by Referee. The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks.
Relatively, Dispatcher and Reasoning Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Reasoning Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity. Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Reasoning Expert have the least impact.
Table 7: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. The red text highlights the performance improvements brought about by the introduction of GPT-4o-mini.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
MME 84.7 (+13.2) 77.8 (+6.3) 71.5 58.8 61.5
MMMU 54.5 (+2.0) 53.4 (+0.9) 52.5 41.5 41.8
4.4 Which One is More Crucial: Visual Understanding or Textual Reasoning?
ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs. However, it remains unclear which of these two capacities is more critical for visual reasoning tasks. To answer this question, we conduct comparative experiments of the following three scenarios:
  1. Llama3-LLaVA-NeXT-8B as All Sub-Agents. All sub-agents within ProReason framework are performed by Llama3-LLaVA-NeXT-8B model.
  2. GPT-4o-mini as Vision Expert. Based on the above scenario, we implement the Vision Expert with GPT-4o-mini, while keep the other textual sub-agents unchanged.
  3. GPT-4o-mini as Textual Sub-Agents. Reversely, we utilize Llama3-LLaVA-NeXT- 8B as the Vision Expert, and GPT-4o-mini for the other vision-irrelevant sub-agents.
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec. 4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
5 Related work
Large Visual-Language Model.
Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c; b; 2024a; Li et al., 2024a; b) and InternVL (Chen et al., 2024b) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Ghosh et al., 2024) and motivating more effective solutions.
Passive Visual Reasoning.
Extracting information from images into text can effectively assist LVLM in performing visual reasoning tasks. Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) first describes the image before prefixing this description to the prompt, assisting LVLMs on visual reasoning tasks. Furthermore, Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) directs LVLMs to generate scene graphs (SGs) that serve as a bridge between the visual and textual domains, aiding LVLMs in subsequent tasks. However, most of these methods employ a question-agnostic and reasoning-free visual extraction process, where image descriptions are not tailored to specific questions, and no reasoning is applied to infer additional information for improved descriptions. These “passive” approaches lead to the inclusion of irrelevant or redundant information, ultimately degrading performance. In contrast, ProReason adopts question-oriented agents to collect necessary and sufficient information, effectively circumventing these drawbacks.
Multi-step Reasoning Framework.
Multi-step reasoning frameworks have been developed for LLMs to achieve better performance by breaking down complex questions into easier ones (Pan et al., 2024). As a representative method, Chain-of-Thought (CoT) (Wei et al., 2022) enhances the arithmetic and commonsense reasoning capabilities by explicitly generating intermediate reasoning steps before concluding the final answers. Tree-of-Thoughts (ToT) (Yao et al., 2024) further refines the CoT mechanism by allowing LLMs to consider multiple reasoning paths and do self-assessment before making decisions. Considering that the inherent knowledge of LLMs may not be sufficient to complete tasks, ReAct (Yao et al., 2022) integrates information retrieval into the reasoning chains, enabling models to pause to verify results and determine whether additional information is needed before proceeding. Nevertheless, multi-step reasoning frameworks designed for text-only LLMs are not fully applicable to visual reasoning tasks, and may even impair the performance of LVLMs (Ghosh et al., 2024).
6 Conclusion
In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities. To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring multi-run proactive perception and decoupled vision-reasoning capabilities. Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models. Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
Appendix A Appendix
A.1 Demonstrative Examples
Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason. ProReason enables LVLMs to proactively acquire necessary information in a question-oriented manner, and predicts answers based on the collected information. Apparently, ProReason is superior to previous methods, which often describe question-irrelevant visual details, or overlook informative elements. Green indicates correct information or conclusions, while red signifies incorrect ones.
Figure 3: Additional examples of images-unrelated Chain-of-Thought reasoning.
Figure 4: Additional examples of question-unrelated passive image captions.
A.2 Prompts
Figure 5: Prompt templates of Relevance Score and caption effectiveness evaluation.
Figure 6: Prompt templates of Chain-of-Thought, fine-grained image captions generation, Compositional Chain-of-Thought (CCoT), and ReAct.
Figure 7: Prompt templates of Dispatcher.
Figure 8: Prompt templates of Vision Expert, Reasoning Expert, Referee, and Summarizer.
A.3 Supplementary Results
Table 8: Average token consumption of ProReason and ReAct with GPT-4o-mini model on the MME and MathVista benchmarks.
Method MME MathVista
Input Output Input Output
ProReason 1286.8 327.2 2238.6 788.6
ReACT 1645.0 197.0 3092.8 845.1
Table 9: Performance of multiple approaches with two base models across four visual reasoning benchmarks. Different from Table 4, MME scores are calculated with the official method here.
Model Method Dataset
MME MMMU MathVista HallusionBench
Llama3-LLaVA- NeXT-8B Direct 359 41.8 37.1 45.8
VDGD 374 42.3 36.1 44.2
CCoT 354 40.5 36.8 37.4
CoT 258 41.5 35.9 43.1
ReAct 385 46.7 31.7 43.6
ProReason 416 52.5 38.8 50.9
GPT-4o-mini Direct 552 48.4 53.0 56.0
VDGD 584 51.4 51.2 52.4
CCoT 535 54.2 53.6 56.7
CoT 729 58.5 53.8 56.3
ReAct 712 54.8 49.3 51.1
ProReason 779 61.6 54.9 58.9
Table 10: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.Different from Table 5, MME scores are calculated with the official method here.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B ReAct 385 46.7 31.7 43.6
ProReason 416 52.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 428 50.4 34.6 40.4
ProReason 603 56.8 48.8 52.3
GPT-4o-mini ReAct 473 48.4 36.2 46.7
ProReason 613 54.5 41.7 53.1
Table 11: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. Different from Table 6, MME scores are calculated with the official method here.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Dispatcher Vision Expert Reasoning Expert Referee Summarizer
MME 779 746 613 743 762 641
MMMU 61.6 60.9 54.5 60.2 51.5 51.0
Table 12: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. Different from Table 7, MME scores are calculated with the official method here.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
MME 613 512 416 286 359
MMMU 54.5 53.4 52.5 41.5 41.8

### v2

@@
 However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation.
-To tackle this issue, we first identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities).
-We then decompose visual reasoning process into two stages: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason.
-This framework features multi-run proactive perception and decoupled vision-reasoning capabilities.
+To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions).
+We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason.
+This framework features decoupled vision-reasoning capabilities and multi-run proactive perception.
 Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions.
 Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs.
-Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on a wide range of benchmarks for both open-source and closed-source models.
-In addition, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark.
-Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
+Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on various benchmarks for both open-source and closed-source models.
+More impressively, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark.
+Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones [^1] Code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
+ProReason: Multi-Modal Proactive Reasoning
+with Decoupled Eyesight and Wisdom
+Jingqi Zhou, Sheng Wang, Jingwei Dong, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
 1 Introduction
@@
 This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks.
-However, state-of-the-art (SOTA) LVLMs still struggle to integrate visual understanding with textual reasoning simultaneously due to inherent modality differences and deficient training data.
+However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences.
 For example, Ghosh et al.
 (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation.
-Figure 2 illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
+Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
 To address the challenges, a promising solution is to extract visual information from images into textual form to assist LVLMs in reasoning (Yin et al., 2023b; Mitra et al., 2024).
@@
 Hence, both of these limitations hinder the performance of LVLMs in multi-modal reasoning tasks.
-To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features proactive (i.e., question-oriented and reasoning-involved) visual extraction and disentangled vision-reasoning capabilities.
-Specifically, we first decouple multi-modal reasoning capacity into two sub-tasks: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom).
-The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions.
-As illustrated in Figure 1, our framework can be divided into three stages: Action, Judgment, and Summary, with each involving different agents.
-In the Action stage, a Dispatcher selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive more insights.
-Unlike passive methods, all three agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency.
-Subsequently, a Referee in the Judgment stage evaluates the sufficiency of known information, deciding whether to return to the Action stage or proceed to the Summary stage, where all information is consolidated for a Summarizer to generate the final answer.
-Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby addressing the limitations of LVLMs.
+To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features disentangled vision-reasoning capabilities and proactive (i.e., question-oriented and reasoning-involved) visual extraction.
+As illustrated in Figure 1, we first decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), based on the capacity of LVLMs.
+The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions, with each involving different agents.
+In the proactive visual perception stage, a Dispatcher decomposes the original question into visual or reasoning-related sub-questions and selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive extra insights.
+Subsequently, a Referee evaluates the sufficiency of known information, deciding whether to continue the proactive visual perception stage or proceed to the textual reasoning stage, where all information is consolidated for a Summarizer to generate the final answer.
+Unlike passive methods, all sub-agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency.
+Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs.
 Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models.
-Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2%.
-Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent, as well as the advantages of proactive information acquisition.
-With the verified superiority over the simultaneous inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
+Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2% (Sec.
+4.2).
+Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent (Sec.
+4.3), as well as the crucial role of decoupling (Sec.
+4.4).
+With the verified superiority over the simultaneously inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks (Sec.
+A.3.1), but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
 The main contributions of this work are threefold:
-  1. We propose a novel multi-modal reasoning framework named ProReason, featuring iterative proactive perception and decoupled vision-reasoning capabilities.
+  1. We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities.
   2.
-Empirically, we identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
+Empirically, we identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
   3. ProReason showcases the remarkable feasibility of integrating LLMs for boosted visual reasoning, illuminating the potential for LLM-assisted LVLM reasoning in future research.
@@
 We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities.
-All experiments are conducted with three recent LVLMs for robustness: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023).
+All experiments are conducted with three recent LVLMs for the robustness of the results: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023).
 Further details are provided in Sec.
@@
 (2024) that LVLMs often rely on prior language knowledge yet neglecting visual information, we hypothesize that the degraded performance might be caused by the image-irrelevant reasoning of CoT.
-To validate the hypothesis, we design a pipeline referencing Liu et al.
+To validate the hypothesis, we design a pipeline inspired by Liu et al.
 (2023c) to measure the correlation between model responses and image contents.
@@
 This highlights the limitation of CoT for multi-modal reasoning tasks, that image-irrelevant reasoning process provokes performance degradation, since the limited capabilities of LVLMs cannot effectively handle both image and text information simultaneously.
-Table 1: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
-Method Model
-Llama3-LLaVA-NeXT-8B LLaVA-OneVision-Qwen2-7B-OV Qwen-VL-Chat
-Direct 41.8 49.1 36.2
-CoT 41.5 46.5 35.4
-VDGD 42.3 49.7 37.1
-Table 2: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
-Model Image Relevance Score
-True False
-Llama3-LLaVA-NeXT-8B 3.70 3.36
-LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
-Qwen-VL-Chat 3.63 2.80
-Table 3: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion.
-“True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
-Score Llama3-LLaVA-NeXT-8B
-True False
-Detail Level 4.43 3.93
-Question Relevance 3.87 3.30
-Reasoning Effective Info Inclusion 3.91 3.57
 2.2 Drawbacks of Passive Information Extraction
@@
 This underscores the limited utility of captions generated by passive methods.
-For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information necessary for reasoning, respectively.
+For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information that is essential for reasoning, respectively.
 Similar to Sec.
@@
 In summary, our analysis highlights the drawbacks of passive visual reasoning enhancement techniques in terms of information insufficiency and redundancy, due to their question-agnostic property.
+Figure 1: Overview and comparison of ProReason, VDGD and ReAct.
+Unlike existing works (e.g., VDGD and ReAct), our proposed method decouples visual perception and textual reasoning while allowing the model to actively acquire necessary information from the images, achieving superior performance.
 (1) ProReason
@@
 (3) Multi-step Reasoning Framework for LLMs
-Figure 1: Overview and comparison of workflows of ProReason, VDGD and ReAct.
-Unlike existing works (e.g., VDGD and ReAct), our proposed method enables the model to actively acquire necessary information from the images, making it more suitable for LVLMs.
 3 Method
-As illustrated in Figure 1, ProReason consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component.
-The entire workflow comprises three steps: Action, Judgment, and Summary, effectively integrating the proactive visual understanding (i.e., eyesight) and textual reasoning (i.e., wisdom) capacities for enhanced multi-modal reasoning performance.
-3.1 Action Step
-Action step is the core of question-oriented visual information extraction, driven by three sub-agents: Dispatcher, Vision Expert, and Reasoning Expert.
-The Dispatcher orchestrates the workflow, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more.
-The responses from both the Vision Expert and the Reasoning Expert are stored in a textual Memory component.
+As depicted in Figure 1, ProReason presents an innovative decoupling of the visual reasoning process into two distinct phases framed through the lens of LVLM capabilities: Proactive Visual Perception (i.e., eyesight) and Textual Reasoning (i.e., wisdom).
+The entire workflow consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component, facilitating multi-modal reasoning performance.
+3.1 Proactive Visual Perception
+Proactive Visual Perception is the core of question-oriented visual information extraction, driven by four sub-agents: Dispatcher, Vision Expert, Reasoning Expert and Referee.
+The Dispatcher breaks down the original question, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more.
+This results in the decoupling of capabilities within the LVLMs, with the Vision Expert handling visual perception and the Reasoning Expert addressing textual reasoning.
+The responses from both experts are stored in a textual Memory component.
+The Referee then evaluates whether the information stored in Memory is sufficient to answer the original question.
 Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Reasoning Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty).
@@
 When the Reasoning Expert is selected, it provides a response consisting of the reasoning process and the final answer $A_{r}$ based on the query $q$ and known information in the Memory, before only $A_{r}$ is stored in Memory.
-Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT and ReAct, thereby suffering less from redundant information.
-3.2 Judgment Step
-Judgment step is conducted by a sub-agent called Referee, which evaluates whether the information stored in Memory is sufficient to answer the question $Q$.
-The input to the Referee includes the question $Q$ and available information in the Memory.
+The Referee then evaluates the available information in the Memory concerning the question $Q$.
 If the Memory contains adequate information to answer the question $Q$, the Referee outputs the identifier “SOLVABLE”; otherwise, it outputs “UNSOLVABLE”.
-If the Referee’s output is SOLVABLE, the workflow precedes to the Summary phase.
-Conversely, if the output is UNSOLVABLE, the Action phase is re-executed to gather more necessary information.
-This allows the Referee to collaborate closely with the three sub-agents in the Action step, enabling the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods.
-3.3 Summary Step
-The Summary step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$.
+If the Referee’s output is SOLVABLE, the workflow precedes to the Textual Reasoning phase.
+Conversely, if the output is UNSOLVABLE, the above process will be re-executed to gather more necessary information.
+In the Proactive Visual Perception phase, four sub-agents collaborate closely to breakdown visual reasoning tasks, thereby achieving an initial decoupling of LVLMs’ visual and reasoning capabilities.
+Simultaneously, this approach enables the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods.
+Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT (Wei et al., 2022) and ReAct (Yao et al., 2022), thereby suffering less from redundant information (for a detailed analysis, refer to Sec.
+A.3.1 and Table 13).
+3.2 Textual Reasoning Step
+The Textual Reasoning step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$.
 This step is mainly powered by a sub-agent called Summarizer.
@@
 3.3 Advantages of ProReason
+Decoupled Visual Reasoning.
+In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents.
+These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
+LLM-assisted multi-modal reasoning.
+On the other hand, this decomposition-before-integration approach not only circumvents the inherent differences between modalities but also allows visually irrelevant sub-agents to be performed by text-only LLMs, thereby facilitating LLM-assisted multi-modal reasoning.
+Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power the multi-modal reasoning process for performance enhancement.
 Reduced information Mission or Redundancy.
-With the close collaboration of the Dispatcher, Vision Expert, Reasoning Expert, and Referee agents, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy.
+With the close collaboration of the sub-agents in Proactive Visual Perception step, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy.
 Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions.
 This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
-LLM-assisted Multi-modal Reasoning.
-In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents.
-These agents are then effectively organized through a designate pipeline to complete the tasks.
-This decomposition-before-integration approach not only circumvents the inherent differences between modalities, but also allows visually irrelevant sub-agents to be performed by text-only LLMs.
-Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power multi-modal reasoning process for performance enhancement.
 4 Experiments
@@
 Datasets.
-To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a).
-All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly as follows:
-  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities.
-Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
-[^2] To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers.
-For the results using the official MME calculation method, please refer to Table 9, Table 10, Table 11, and Table 12.
-  • MMMU evaluates multi-modal models with multidisciplinary tasks that require college-level domain-specific knowledge and detailed reasoning.
-It comprises 11,500 questions across 30 disciplines and 183 sub-fields, emphasizing advanced perception and domain-specific reasoning.
-  • MathVista focuses on more challenging mathematical reasoning tasks that demand precise visual recognition and compositional reasoning.
-It includes 6,141 examples from 31 multi-modal mathematics datasets.
-  • HallusionBench evaluates models’ ability to reason with images such as statistical charts, emphasizing nuanced visual understanding.
-It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
+To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a)  [^3] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
+To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers.
+For the results using the official MME calculation method, please refer to Table 16, Table 16, Table 18, Table 16 and Table 18.
+, Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a).
+All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Sec.
+A.2.
 Base Models.
-We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources.
+We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B [^4] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards.
+and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources.
 As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API.
@@
 Implementation Details.
-The prompt templates for all methods are shown in Figures 6, 7 and 8.
-Specifically for ProReason, in order to prevent an infinite loop, if Dispatcher selects Vision Expert or Reasoning Expert to obtain information up to 5 times, and Referee still judges that the existing knowledge in Memory is not enough to solve the question, the information collection phase will be terminated immediately, and then Summarizer will directly answer the question based on the available information in Memory.
+The prompt templates for all methods are shown in Figures 9, 10 and 12.
+Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Reasoning Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process.
+If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated.
+Subsequently, the Summarizer will generate the final answer based on the available information in Memory from the last attempted iteration.
 This setup, refined through multiple trials, is the most effective.
@@
 ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
-ProReason 71.5 (+10.0) 52.5 (+10.7) 38.8 (+1.7) 50.9 (+5.1)
-Average 66.30 44.22 36.06 44.16
+ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1)
+Average 66.30 43.9 36.06 44.16
 GPT-4o-mini Direct 79.2 48.4 53.0 56.0
@@
 ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
-ProReason 90.4 (+11.2) 61.6 (+13.2) 54.9 (+1.9) 58.9 (+2.9)
-Average 84.63 54.82 52.63 55.23
+ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9)
+Average 84.63 54.82 52.63 55.4
 Table 2: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.
@@
 ReAct 68.5 46.7 31.7 43.6
-ProReason 71.5 52.5 38.8 50.9
+ProReason 71.5 50.5 38.8 50.9
 Qwen2.5-72B- Instruct ReAct 71.0 (+2.5) 50.4 (+3.7) 34.6 (+2.9) 40.4 (-3.2)
-ProReason 81.3 (+9.8) 56.8 (+4.3) 48.8 (+10.0) 52.3 (+1.4)
+ProReason 81.3 (+9.8) 56.8 (+6.3) 48.8 (+10.0) 52.3 (+1.4)
 GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
-ProReason 84.7 (+13.2) 54.5 (+2.0) 41.7 (+2.9) 53.1 (+2.2)
+ProReason 84.7 (+13.2) 54.5 (+4.0) 41.7 (+2.9) 53.1 (+2.2)
 4.2 Main Results
@@
 As mentioned in Sec.
-3.4, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs.
+3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs.
 To demonstrate the utility of this advantage, we fix the Vision Expert as Llama3-LLaVA-NeXT-8B, and replace other agents with text-only LLMs.
@@
 The experimental results are presented in Table 13.
-Table 6: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
-For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
-The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
-Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
-Dispatcher Vision Expert Reasoning Expert Referee Summarizer
-MME 90.4 88.8 (-1.6) 84.7 (-5.7) 88.7 (-1.7) 89.6 (-1.1) 84.2 (-6.2)
-MMMU 61.6 60.9 (-0.7) 54.5 (-7.1) 60.2 (-1.4) 51.5 (-10.1) 51.0 (-10.6)
 Summarizer is the most crucial sub-agent, closely followed by Referee.
@@
 Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks.
+The analysis in Section A.3.2 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
 Relatively, Dispatcher and Reasoning Expert are the least essential sub-agents.
@@
 Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Reasoning Expert have the least impact.
-Table 7: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks.
-The red text highlights the performance improvements brought about by the introduction of GPT-4o-mini.
-Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
-Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
-MME 84.7 (+13.2) 77.8 (+6.3) 71.5 58.8 61.5
-MMMU 54.5 (+2.0) 53.4 (+0.9) 52.5 41.5 41.8
-4.4 Which One is More Crucial: Visual Understanding or Textual Reasoning?
-ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs.
-However, it remains unclear which of these two capacities is more critical for visual reasoning tasks.
-To answer this question, we conduct comparative experiments of the following three scenarios:
-  1. Llama3-LLaVA-NeXT-8B as All Sub-Agents. All sub-agents within ProReason framework are performed by Llama3-LLaVA-NeXT-8B model.
-  2. GPT-4o-mini as Vision Expert. Based on the above scenario, we implement the Vision Expert with GPT-4o-mini, while keep the other textual sub-agents unchanged.
-  3. GPT-4o-mini as Textual Sub-Agents. Reversely, we utilize Llama3-LLaVA-NeXT- 8B as the Vision Expert, and GPT-4o-mini for the other vision-irrelevant sub-agents.
-Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important.
-As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities.
-However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert.
-This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec.
-4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
-5 Related work
+4.4 The Critical Implication of Decoupling
+To validate the critical implication of decoupling visual perception and textual reasoning in ProReason while eliminating potential confounding factors from prompt engineering and multiple CoT implementations, we systematically integrate sub-agents through three configurations:
+  1. Merge the Vision Expert and Reasoning Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
+  2. Integrate the Dispatcher, Vision Expert, Reasoning Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
+  3. Fully consolidate all five original sub-agents to demonstrate the pivotal role of decomposing visual reasoning tasks into distinct Proactive Visual Perception and Textual Reasoning phases.
+The merged agents preserve identical prompts and maintain the same task execution procedures as their original counterparts, with their prompts shown in Figures 13, 14, and 15.
+Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
+Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks.
+As shown in Table 13, merging Vision and Reasoning Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs.
+CoT’s 58.5).
+Although agent merging also causes performance declines on MME, the merged versions still outperform CoT.
+Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
+5 Conclusion
+In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities.
+To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring decoupled vision-reasoning capabilities and multi-run proactive perception.
+Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models.
+Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
+6 Limitations
+In this section, we analyze the limitations of the proposed method based on typical errors made by GPT-4o-mini-driven ProReason, as exemplified in Figures 6 and 7, to gain further understanding and identify potential research directions.
+Cumulative Errors.
+As illustrated in Figure 6, the vision expert mistakenly perceives the clock as 6:25, which misguides the reasoning of subsequent agents and ultimately leads to an incorrect conclusion.
+More broadly, similar misperceptions occur frequently in errors made by ProReason.
+This indicates that, with the assistance of LLMs, ProReason has effectively addressed the reasoning deficiencies in multi-modal tasks, while the vision expert plays a significant role for further improvement of multi-modal capabilities.
+Contradictory Information among Agents.
+Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and reasoning expert.
+However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the reasoning expert) tend to adhere to the available information instead of questioning it.
+This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
+7 Ethics Statement
+We adhere strictly to the ACL Code of Ethics throughout our research.
+To our knowledge, the methods we introduce pose no foreseeable risks.
+We provide comprehensive details of the computing infrastructure used for all computational experiments in the paper, along with transparent statistics on our results and a detailed configuration of our experimental setup, including the optimal hyperparameter values.
+Furthermore, we will release the code upon publication to facilitate easy public reproducibility.
+Appendix A Appendix
+A.1 Related work
 Large Visual-Language Model.
 Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents.
-In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c; b; 2024a; Li et al., 2024a; b) and InternVL (Chen et al., 2024b) families, have been extensively developed.
+In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024b) families, have been extensively developed.
 In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success.
@@
 Nevertheless, multi-step reasoning frameworks designed for text-only LLMs are not fully applicable to visual reasoning tasks, and may even impair the performance of LVLMs (Ghosh et al., 2024).
-6 Conclusion
-In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities.
-To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring multi-run proactive perception and decoupled vision-reasoning capabilities.
-Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models.
-Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
-Appendix A Appendix
-A.1 Demonstrative Examples
-Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason.
-ProReason enables LVLMs to proactively acquire necessary information in a question-oriented manner, and predicts answers based on the collected information.
-Apparently, ProReason is superior to previous methods, which often describe question-irrelevant visual details, or overlook informative elements.
-Green indicates correct information or conclusions, while red signifies incorrect ones.
-Figure 3: Additional examples of images-unrelated Chain-of-Thought reasoning.
-Figure 4: Additional examples of question-unrelated passive image captions.
-A.2 Prompts
-Figure 5: Prompt templates of Relevance Score and caption effectiveness evaluation.
-Figure 6: Prompt templates of Chain-of-Thought, fine-grained image captions generation, Compositional Chain-of-Thought (CCoT), and ReAct.
-Figure 7: Prompt templates of Dispatcher.
-Figure 8: Prompt templates of Vision Expert, Reasoning Expert, Referee, and Summarizer.
-A.3 Supplementary Results
+A.2 Dataset
+To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a).
+Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
+  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities.
+Given our focus on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
+  • MMMU evaluates multi-modal models with multidisciplinary tasks that require college-level domain-specific knowledge and detailed reasoning.
+It comprises 11,500 questions across 30 disciplines and 183 sub-fields, emphasizing advanced perception and domain-specific reasoning.
+  • MathVista focuses on more challenging mathematical reasoning tasks that demand precise visual recognition and compositional reasoning.
+It includes 6,141 examples from 31 multi-modal mathematics datasets.
+  • HallusionBench evaluates models’ ability to reason with images such as statistical charts, emphasizing nuanced visual understanding.
+It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
+A.3 Supplementary Results and Analysis
+A.3.1 Which One is More Crucial: Visual Understanding or Textual Reasoning?
+ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs.
+However, it remains unclear which of these two capacities is more critical for visual reasoning tasks.
+To answer this question, we conduct comparative experiments of the following three scenarios:
+  1. Llama3-LLaVA-NeXT-8B as All Sub-Agents. All sub-agents within ProReason framework are performed by Llama3-LLaVA-NeXT-8B model.
+  2. GPT-4o-mini as Vision Expert. Based on the above scenario, we implement the Vision Expert with GPT-4o-mini, while keep the other textual sub-agents unchanged.
+  3. GPT-4o-mini as Textual Sub-Agents. Reversely, we utilize Llama3-LLaVA-NeXT- 8B as the Vision Expert, and GPT-4o-mini for the other vision-irrelevant sub-agents.
+Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important.
+As shown in Table 3, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities.
+However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert.
+This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec.
+4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
+Table 3: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks.
+The red text highlights the performance improvements brought about by the introduction of GPT-4o-mini.
+Model Agent Dataset
+MME MMMU
+GPT-4o -mini Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
+Vision Expert 77.8 (+6.3) 53.4 (+0.9)
+Llama3- LLaVA- NeXT-8B All Sub-Agents 71.5 52.5
+COT 58.8 41.5
+Direct 61.5 41.8
+A.3.2 Referee’s Dispel of Hallucinations
+Following the implementation details outlined in Section 4.1, we evaluate the ProReason system powered by GPT-4o-mini on both MMMU and HallusionBench under varying attempt allowances (i.e., 1, 3, and 5 attempts).
+Every unsuccessful attempt reflects the Referee’s persistent determination that the information stored in Memory is insufficient to solve the problem.
+Before each new attempt, Memory is systematically cleared, ensuring the removal of information deemed irrelevant by the Referee’s assessment.
+Consequently, increased attempt allowances essentially empower the Referee with enhanced opportunities for information filtration.
+The observed performance variations of ProReason across different attempt quotas demonstrate the critical impact of Referee’s decision-making mechanism and information filtering efficacy on system capability.
+The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework.
+As demonstrated in Table 4, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5.
+Given HallusionBench’s dual emphasis on reasoning proficiency and precise evaluation of visual details/hallucination control compared to MMMU, these results suggest that the Referee mechanism can effectively identify erroneous or irrelevant visual information, thereby strengthening ProReason’s capacity for meticulous visual understanding.
+Table 4: Impact on HallusionBench and MMMU performance across different attempt allowances.
+Dataset Attempt Allowances
+1 3 5
+MMMU 60.1 59.9 61.6
+HallusionBench 57.1 58.9 59.9
+A.3.3 Performance of ProReason on VQA task
+ProReason also excels in conventional VQA tasks.
+We perform an evaluation on the VQA benchmark A-OKVQA dataset (Schwenk et al., 2022).
+The visual complexity of A-OKVQA is comparable to MME, but its reasoning requirements are less demanding.
+Table 5 illustrates the performance of ProReason, powered by GPT-4o-mini, on the A-OKVQA, MME, and MMMU benchmarks.
+ProReason achieves the highest performance across all three benchmarks, indicating the positive impact of ProReason on any form of reasoning-involved multi-modal tasks.
+ProReason is particularly effective on datasets with greater complexity due to its decoupled capabilities and LLM-assisted reasoning.
+Table 5 demonstrates that ProReason outperforms the "Direct" method by 2.7%, 11.2%, and 13.2% on A-OKVQA, MME, and MMMU, respectively, as the visual and reasoning complexity increases.
+This underscores the advantage of ProReason in handling more challenging tasks, making it a robust choice for complex problem-solving scenarios.
+Table 5: Performance of multiple approaches with GPT-4o-mini across three visual benchmarks.
+GPT-4o-mini Dataset
+A-OKVQA MME MMMU
+Direct 78.9 79.2 48.4
+VDGD 79.4 82.3 51.4
+CCoT 79.2 80.8 54.2
+COT 80.9 87.8 58.5
+ReAct 80.6 87.3 54.8
+ProReason 81.3 (+2.7) 90.4 (+11.2) 61.6 (+13.2)
+A.3.4 Evaluation of ProReason based on LLMs
+In Sec.
+2.1, we utilize a pipeline based on LLMs to evaluate the correlation between model responses and image content.
+We aim to apply this effective method once more to assess the performance of ProReason.
+Initially, we adopted the reasoning chains generated by GPT-4o-mini (CoT) as the standard answers, given its superior performance with a score of 58.5 on the MMMU benchmark, compared to other models.
+However, due to ProReason’s enhancement of GPT-4o-mini’s score to 61.6, the use of GPT-4o-mini’s answers became inadequate.
+Consequently, we opted for GPT-4o to generate the standard answers, achieving a score of 69.1 on MMMU.
+As shown in Table 6, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason (ProReason-Memory), and the reasoning process of Summarizer (ProReason-Summarizer), which generates the final answer of ProReason with the prompt template in Figure 12.
+This assessment focuses on three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement.
+Specifically, compared to GPT-4o-mini (CoT), ProReason-Summarizer produces more relevant answers with less redundancy and deficiency, aligning with its improved performance.
+Compared to ProReason-Summarizer, ProReason-Memory exhibits the same RE, higher RI and lower MI scores.
+This suggests that ProReason allows some redundancy to prevent information loss in memory, as the former typically leads to more serious consequences than the latter.
+Subsequently, Summarizer can leverage its powerful reasoning capabilities to select the most relevant memory.
+Table 6: Performance of ProReason driven by GPT-4o-mini assessed by LLMs compared to CoT on MMMU benchmark.
+Mainly includes three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI ↓), where arrows indicate the directions of improvement.
+GPT-4o-mini Metrics
+RE ↑ RI ↓ MI ↓
+COT 4.67 3.33 1.40
+ProReason-Memory 4.83 3.66 1.17
+ProReason-Summarizer 4.83 2.88 1.33
+A.3.5 Frequency of selection of various experts
+As listed in Table 7, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities.
+Specifically, compared to MME, the frequencies for both the Vision and Reasoning Experts are higher on the MMMU benchmark, aligning with their difficulty levels.
+Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%).
+Despite the lower frequency of the reasoning expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions.
+Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
+Table 7: Frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks.
+Dataset GPT-4o-mini
+Vision Expert Reasoning Expert
+MME 1.16 0.12
+MMMU 1.64 0.38
+Table 8: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
+Model Method
+Direct CoT VDGD
+Llama3-LLaVA-NeXT-8B 41.8 41.5 42.3
+LLaVA-OneVision-Qwen2-7B-OV 49.1 46.5 49.7
+Qwen-VL-Chat 36.2 35.4 37.1
+Table 9: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
+Model Image Relevance Score
+True False
+Llama3-LLaVA-NeXT-8B 3.70 3.36
+LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
+Qwen-VL-Chat 3.63 2.80
+Table 10: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion.
+“True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
+Score Llama3-LLaVA-NeXT-8B
+True False
+Detail Level 4.43 3.93
+Question Relevance 3.87 3.30
+Reasoning Effective Info Inclusion 3.91 3.57
 Table 11: Average token consumption of ProReason and ReAct with GPT-4o-mini model on the MME and MathVista benchmarks.
@@
 ReACT 1645.0 197.0 3092.8 845.1
+Table 12: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
+For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
+The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
+Model Agent Dataset
+MME MMMU
+GPT-4o-mini 90.4 61.6
+Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
+Vision Expert 84.7 (-5.7) 54.5 (-7.1)
+Reasoning Expert 88.7 (-1.7) 60.2 (-1.4)
+Referee 89.6 (-1.1) 51.5 (-10.1)
+Summarizer 84.2 (-6.2) 51.0 (-10.6)
+Table 13: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments.
+The blue text highlights the performance degradation due to sub-agent integration.
+GPT-4o-mini Dataset
+MMMU MME
+ProReason 61.6 91.9
+CoT 58.5 87.8
+Vision & Reasoning
+Expert Integration
+58.6 (-3.0) 90.4 (-1.5)
+Dispatcher,
+Vision & Reasoning Expert,
+& Referee Integration
+57.8 (-3.8) 90.8 (-1.1)
+All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
 Table 14: Performance of multiple approaches with two base models across four visual reasoning benchmarks. Different from Table 1, MME scores are calculated with the official method here.
@@
 MME MMMU MathVista HallusionBench
-Llama3-LLaVA- NeXT-8B Direct 359 41.8 37.1 45.8
-VDGD 374 42.3 36.1 44.2
-CCoT 354 40.5 36.8 37.4
-CoT 258 41.5 35.9 43.1
-ReAct 385 46.7 31.7 43.6
-ProReason 416 52.5 38.8 50.9
-GPT-4o-mini Direct 552 48.4 53.0 56.0
-VDGD 584 51.4 51.2 52.4
-CCoT 535 54.2 53.6 56.7
-CoT 729 58.5 53.8 56.3
-ReAct 712 54.8 49.3 51.1
-ProReason 779 61.6 54.9 58.9
-Table 10: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.Different from Table 5, MME scores are calculated with the official method here.
+Llama3-LLaVA- NeXT-8B Direct 35.9 41.8 37.1 45.8
+VDGD 37.4 42.3 36.1 44.2
+CCoT 35.4 40.5 36.8 37.4
+CoT 25.8 41.5 35.9 43.1
+ReAct 38.5 46.7 31.7 43.6
+ProReason 41.6 50.5 38.8 50.9
+GPT-4o-mini Direct 55.2 48.4 53.0 56.0
+VDGD 58.4 51.4 51.2 52.4
+CCoT 53.5 54.2 53.6 56.7
+CoT 72.9 58.5 53.8 56.3
+ReAct 71.2 54.8 49.3 51.1
+ProReason 77.9 61.6 54.9 59.9
+Table 15: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.
+Different from Table 2, MME scores are calculated with the official method here.
 Vision Expert Textual Sub-Agents Method Dataset
@@
 Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B ReAct 385 46.7 31.7 43.6
-ProReason 416 52.5 38.8 50.9
+ProReason 416 50.5 38.8 50.9
 Qwen2.5-72B- Instruct ReAct 428 50.4 34.6 40.4
@@
 ProReason 613 54.5 41.7 53.1
-Table 11: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
-For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
-Different from Table 6, MME scores are calculated with the official method here.
-Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
-Dispatcher Vision Expert Reasoning Expert Referee Summarizer
-MME 779 746 613 743 762 641
-MMMU 61.6 60.9 54.5 60.2 51.5 51.0
 Table 16: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks.
@@
 MMMU 54.5 53.4 52.5 41.5 41.8
+Table 17: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
+For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
+Different from Table 13, MME scores are calculated with the official method here.
+Model Agent Dataset
+MME MMMU
+GPT-4o-mini 779 61.6
+Llama3- LLaVA- NeXT-8B Dispatcher 746 60.9
+Vision Expert 613 54.5
+Reasoning Expert 743 60.2
+Referee 762 51.5
+Summarizer 641 51.0
+Table 18: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments.
+Different from Table 13, MME scores are calculated with the official method here.
+GPT-4o-mini Dataset
+MMMU MME
+ProReason 61.6 779
+CoT 58.5 729
+Vision & Reasoning
+Expert Integration
+58.6 673
+Dispatcher,
+Vision & Reasoning Expert,
+& Referee Integration
+57.8 714
+All Sub-Agents Integration 56.8 711
+A.4 Demonstrative Examples
+Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason.
+ProReason enables LVLMs to proactively acquire necessary information in a question-oriented manner, and predicts answers based on the collected information.
+Apparently, ProReason is superior to previous methods, which often describe question-irrelevant visual details, or overlook informative elements.
+Green indicates correct information or conclusions, while red signifies incorrect ones.
+Figure 3: A complete reasoning process of ProReason for the case shown in Figure 2.
+Figure 4: Additional examples of images-unrelated Chain-of-Thought reasoning.
+Figure 5: Additional examples of question-unrelated passive image captions.
+Figure 6: A typical mistake made by ProReason.
+The vision expert incorrectly identifies 4:30 as 6:25, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
+Figure 7: A typical mistake made by ProReason.
+The vision expert incorrectly provided the coordinates of the three vertices of the triangle, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
+A.5 Prompts
+Figure 8: Prompt templates of Relevance Score and caption effectiveness evaluation.
+Figure 9: Prompt templates of Chain-of-Thought, fine-grained image captions generation, Compositional Chain-of-Thought (CCoT), and ReAct.
+Figure 10: Prompt templates of Dispatcher.
+Figure 11: Prompt templates of Vision Expert, Reasoning Expert, Referee, and Summarizer.
+Figure 12: Prompt templates of LLMs-based ProReason evaluation.
+Figure 13: Prompt templates of Vision and Reasoning Expert Integration.
+Figure 14: Prompt templates of Dispatcher, Vision Expert Reasoning Expert and Referee Integration.
+Figure 15: Prompt templates of All Five Sub-Agent Integration.

### v3

@@
 Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs.
-Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on various benchmarks for both open-source and closed-source models.
-More impressively, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark.
-Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones [^1] Code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
+Our extensive experiments demonstrate that ProReason outperforms existing multi-step reasoning frameworks on various benchmarks for both open-source and closed-source models, with the average performance gain reaching 13.2%.
+Besides, the integration of LLMs allows ProReason to produce high-quality visual reasoning data, which empowers ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) to achieve superior performance in downstream tasks.
+Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
 ProReason: Multi-Modal Proactive Reasoning
 with Decoupled Eyesight and Wisdom
-Jingqi Zhou, Sheng Wang, Jingwei Dong, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
+Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Lui, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, u3638070@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
 1 Introduction
-In recent years, large language models (LLMs) (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024).
-This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks.
+In recent years, large language models (LLMs) (Yang et al., 2024; Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024).
+This progress has also sparked interest in developing large vision-language models (LVLMs) (Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks.
 However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences.
 For example, Ghosh et al.
-(2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation.
+(2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks, such as visual chart understanding and math reasoning, resulting in performance degradation.
 Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
-To address the challenges, a promising solution is to extract visual information from images into textual form to assist LVLMs in reasoning (Yin et al., 2023b; Mitra et al., 2024).
-For example, Ghosh et al.
-(2024) instruct LVLMs to generate fine-grained captions to facilitate the subsequent reasoning process.
-However, these existing methods exhibit two primary limitations: (i) The visual extraction process is question-agnostic and reasoning-free, that the image description is not targeted for a given question, and reasoning is not involved to deduce extra information for better descriptions.
-This drawback, termed as “passive”, results in irrelevant or insufficient information, and subsequent performance degradation.
-(ii) The reasoning process entangles visual understanding with textual reasoning abilities of a single LVLM.
-As mentioned above, the limited capabilities of LVLMs do not always successfully handle both abilities and produce high-quality reasoning.
-Hence, both of these limitations hinder the performance of LVLMs in multi-modal reasoning tasks.
-To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features disentangled vision-reasoning capabilities and proactive (i.e., question-oriented and reasoning-involved) visual extraction.
-As illustrated in Figure 1, we first decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), based on the capacity of LVLMs.
-The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions, with each involving different agents.
-In the proactive visual perception stage, a Dispatcher decomposes the original question into visual or reasoning-related sub-questions and selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive extra insights.
-Subsequently, a Referee evaluates the sufficiency of known information, deciding whether to continue the proactive visual perception stage or proceed to the textual reasoning stage, where all information is consolidated for a Summarizer to generate the final answer.
-Unlike passive methods, all sub-agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency.
-Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs.
-Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models.
-Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2% (Sec.
-4.2).
-Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent (Sec.
-4.3), as well as the crucial role of decoupling (Sec.
-4.4).
-With the verified superiority over the simultaneously inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks (Sec.
-A.3.1), but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
+To address the challenges, several visual reasoning frameworks have been proposed.
+Specifically, Ghosh et al.
+(2024) and Mitra et al.
+(2024) convert visual information in images into textual descriptions to aid LVLMs in reasoning.
+However, their visual extraction process is not targeted at a given question (i.e., question-agnostic), termed as "passive", and omits reasoning mechanisms to infer extra information for better descriptions(i.e., reasoning-free).
+These limitations result in irrelevant or inadequate information, ultimately degrading performance.
+Furthermore, these frameworks are powered by a single LVLM, leading to a reasoning process that conflates visual understanding with textual reasoning abilities, failing to mitigate the challenge faced by LVLMs in effectively managing both capabilities.
+To resolve these problems, we propose ProReason, a multi-modal reasoning framework featuring decoupled vision-reasoning capabilities.
+As illustrated in Figure 1, we decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom).
+The former extracts visual information in a question-oriented and reasoning-involved manner, while the latter integrates all information to draw final conclusions.
+Specifically, during the visual perception stage, a Dispatcher first selectively engages a Vision Expert to capture additional visual information, or an Insight Expert to derive intermediate inferences.
+A Referee then determines whether sufficient information is gathered to proceed to the reasoning stage, where a Summarizer produces the final answer.
+Unlike passive methods, all sub-agents operate based on the given question and known information, effectively avoiding irrelevant information redundancy or insufficiency.
+Notably, decoupled vision-reasoning eliminates the need for LVLMs to handle vision-irrelevant roles (i.e., Dispatcher, Insight Expert, Referee, and Summarizer), enabling seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs.
+In addition, the high-quality reasoning data generated by LLM-assisted ProReason can be effectively distilled into downstream models for inherent performance improvement.
+Empirically, we evaluate ProReason across multiple challenging visual reasoning benchmarks with both open-source and closed-source models.
+Extensive experiments demonstrate that ProReason exhibits significant advantages in two key aspects: (1) As a visual reasoning framework, ProReason achieves consistent and substantial performance improvements across multiple benchmarks, with the average performance gain reaching 13.2%, validating the effectiveness of its decoupled vision-reasoning architecture and proactive visual feature extraction mechanism; (2) ProReason effectively integrates existing LLMs to generate high-quality visual reasoning process, empowering ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) with superior visual reasoning capabilities.
+The above results, coupled with the ablation study in Section 4.5, demonstrate the substantial advantages of decoupled vision-reasoning, while highlighting the potential of LLM-assisted LVLM reasoning and distillation strategies.
 The main contributions of this work are threefold:
-  1. We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities.
-  2.
-Empirically, we identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
-  3. ProReason showcases the remarkable feasibility of integrating LLMs for boosted visual reasoning, illuminating the potential for LLM-assisted LVLM reasoning in future research.
+  • We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities, effectively mitigating the drawbacks of previous methods.
+  • Extensive experiments consistently highlight the significant superiority of ProReason and necessity of each component across multiple visual reasoning tasks and model series, illuminating the great potential of LLM-assisted LVLM reasoning.
+  • ProReason-distilled models also exhibits remarkable enhancements over vanilla counterparts, showcasing the feasibility of LLM-assisted LVLM improvement in the future.
 2 Preliminary Observations
-We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities.
-All experiments are conducted with three recent LVLMs for the robustness of the results: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023).
-Further details are provided in Sec.
-4.1.
-2.1 Challenges in Visual Reasoning: Limitations of LVLMs
-Chain-of-Thought(CoT) (Wei et al., 2022) has been extensively verified to enhance the performance of LLMs.
-Here, we explore its impact on the reasoning performance of LVLMs.
-Similarly, we require the models to “think step by step” before generating final answers.
-As shown in Table 8, a counter-intuitive phenomenon is observed: compared to “Direct” answering method, the introduction of CoT consistently incurs slight performance degradation across all three models.
-Inspired by the findings of Ghosh et al.
-(2024) that LVLMs often rely on prior language knowledge yet neglecting visual information, we hypothesize that the degraded performance might be caused by the image-irrelevant reasoning of CoT.
-To validate the hypothesis, we design a pipeline inspired by Liu et al.
-(2023c) to measure the correlation between model responses and image contents.
-Specifically, given a question, we take the reasoning process generated by GPT-4o-mini [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence, whose superiority on MMMU task is verified in Sec.
-4.2, as the golden solution.
-Subsequently, a more powerful judge (i.e., GPT-4 (Achiam et al., 2023)) is employed to evaluate the relevance of responses of candidate models (e.g., Llama3-LLaVA-NeXT-8B) and this golden solution with the prompt template in Figure 8, and gives a comprehensive score from 1 to 5, with 5 indicating the highest relevance.
-Since the golden solution contains the essential information needed to solve the problem within the image, a higher correlation indicates a stronger relevance to the image content.
-As shown in Table 9, the reasoning chains of incorrect responses are significantly less relevant to the images than those of correct answers.
-An exemplary case is shown in Figure 2, where Llama3-LLaVA-NeXT-8B attempts to use the right-hand rule to solve the task, but its reasoning process is almost unrelated to the image, ultimately leading to an incorrect conclusion.
-This highlights the limitation of CoT for multi-modal reasoning tasks, that image-irrelevant reasoning process provokes performance degradation, since the limited capabilities of LVLMs cannot effectively handle both image and text information simultaneously.
-2.2 Drawbacks of Passive Information Extraction
-Passive visual reasoning techniques suffer insufficient and irrelevant visual information, despite mitigating the oversight of images by converting them into detailed captions, like Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024).
-To support this claim, we generate fine-grained image captions using GPT-4o-mini with the prompt shown in Figure 9.
+Ghosh et al.
+(2024) demonstrate that the limited multimodal reasoning abilities of LVLMs lead to an overreliance on linguistic priors, thus neglecting visual inputs and ultimately degrading their performance.
+Their Visual Description Grounded Decoding (VDGD) mitigates visual oversight by converting images into comprehensive textual descriptions to inform reasoning processes.
+However, such passive visual reasoning techniques suffer insufficient and irrelevant visual information.
+To support this claim, we generate fine-grained image captions using GPT-4o-mini [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence with the prompt shown in Figure 9.
 We then incorporate these captions into the prompts for LVLMs to facilitate the reasoning process.
-As listed in Table 8, while these image descriptions improve the performance of LVLMs, the gains are marginal, consistently amounting to less than 1%.
+We analyze the performance of this approach on the challenging multi-modal MMMU dataset, which requires college-level knowledge and fine-grained reasoning, using recent open-source LVLMs listed in Section 4.1 As shown in Table 4, while these image descriptions improve the performance of LVLMs, the gains are marginal, consistently amounting to less than 1%.
 This underscores the limited utility of captions generated by passive methods.
-For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information that is essential for reasoning, respectively.
-Similar to Sec.
-2.1, we pair the image captions with the reasoning process of GPT-4o-mini, and employ GPT-4 as a judge to rate the captions from 1 to 5 across these dimensions.
-The prompt template is shown in Figure 8.
+For further demonstration, inspired by Liu et al.
+(2023c), we use the prompt instructions in Figure 8 to instruct GPT-4 to analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information that is essential for reasoning, respectively.
+Meanwhile, since the reasoning process of GPT-4o-mini on MMMU contains key information necessary for solving the problems, we use it as a reference answer to aid evaluation.
 As shown in Table 5, the captions for correct responses of Llama3-LLaVA-NeXT-8B receive higher scores across all three criteria, highlighting the importance of better captions for multi-modal reasoning.
@@
 3 Method
-As depicted in Figure 1, ProReason presents an innovative decoupling of the visual reasoning process into two distinct phases framed through the lens of LVLM capabilities: Proactive Visual Perception (i.e., eyesight) and Textual Reasoning (i.e., wisdom).
+As depicted in Figure 1 and exemplified in Figure 3, ProReason presents an innovative decoupling of the visual reasoning process into two distinct phases framed through the lens of LVLM capabilities: Proactive Visual Perception (i.e., eyesight) and Textual Reasoning (i.e., wisdom).
 The entire workflow consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component, facilitating multi-modal reasoning performance.
 3.1 Proactive Visual Perception
-Proactive Visual Perception is the core of question-oriented visual information extraction, driven by four sub-agents: Dispatcher, Vision Expert, Reasoning Expert and Referee.
-The Dispatcher breaks down the original question, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more.
-This results in the decoupling of capabilities within the LVLMs, with the Vision Expert handling visual perception and the Reasoning Expert addressing textual reasoning.
+Proactive Visual Perception is the core of question-oriented visual information extraction, driven by four sub-agents: Dispatcher, Vision Expert, Insight Expert and Referee.
+The Dispatcher breaks down the original question, selectively directing the Vision Expert to capture specific visual information, or instructing the Insight Expert to analyze known information to derive more.
 The responses from both experts are stored in a textual Memory component.
 The Referee then evaluates whether the information stored in Memory is sufficient to answer the original question.
-Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Reasoning Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty).
+Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Insight Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty).
 The Dispatcher then generates a query $q$ for the chosen expert.
-If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{v}$, which is then stored in the Memory.
-When the Reasoning Expert is selected, it provides a response consisting of the reasoning process and the final answer $A_{r}$ based on the query $q$ and known information in the Memory, before only $A_{r}$ is stored in Memory.
+If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{vision}$, which is then stored in the Memory.
+When the Insight Expert is selected, it provides a response consisting of the inference process and the final answer $A_{insight}$ based on the query $q$ and known information in the Memory, before only $A_{insight}$ is stored in Memory.
 The Referee then evaluates the available information in the Memory concerning the question $Q$.
@@
 Conversely, if the output is UNSOLVABLE, the above process will be re-executed to gather more necessary information.
-In the Proactive Visual Perception phase, four sub-agents collaborate closely to breakdown visual reasoning tasks, thereby achieving an initial decoupling of LVLMs’ visual and reasoning capabilities.
-Simultaneously, this approach enables the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods.
-Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT (Wei et al., 2022) and ReAct (Yao et al., 2022), thereby suffering less from redundant information (for a detailed analysis, refer to Sec.
-A.3.1 and Table 13).
+In the Proactive Visual Perception phase, the Dispatcher collaborates with the Vision Expert and Insight Expert to achieve question-oriented and reasoning-involved visual information extraction, while the Referee ensures informational completeness and prevents omissions.
+These four sub-agents work closely together, thereby overcoming the limitations of passive methods.
+Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT (Wei et al., 2022) and ReAct (Yao et al., 2022), thereby suffering less from redundant information (for a detailed analysis, refer to Section 4.4 and Table 3).
 3.2 Textual Reasoning Step
@@
 This step is mainly powered by a sub-agent called Summarizer.
-Once the Referee determines that the information in the Memory is sufficient to address the question $Q$, and outputs the identifier SOLVABLE, the Summarizer will be called to draw a conclusion, based on the Memory information.
-This conclusion represents the answer to the question $Q$ of the whole framework, and will be evaluated by the performance metrics.
+Once the Referee confirms that the Memory contains sufficient information to address question $Q$ and outputs the SOLVABLE identifier, the Summarizer will be invoked to perform a detailed reasoning and generate a final answer based on the question $Q$ and Memory.
+This final answer is then evaluated using performance metrics, while the Summarizer’s high-quality reasoning outputs can be utilized to train downstream models, enhancing their visual reasoning capabilities.
 3.3 Advantages of ProReason
 Decoupled Visual Reasoning.
-In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents.
+In ProReason, the multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents.
 These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
 LLM-assisted multi-modal reasoning.
-On the other hand, this decomposition-before-integration approach not only circumvents the inherent differences between modalities but also allows visually irrelevant sub-agents to be performed by text-only LLMs, thereby facilitating LLM-assisted multi-modal reasoning.
-Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power the multi-modal reasoning process for performance enhancement.
+Decoupled vision-reasoning eliminates the necessity for LVLMs to manage vision-irrelevant sub-agents, allowing the seamless integration of existing LLMs with established strong reasoning abilities, thereby endowing ProReason with superior visual reasoning performance.
+Downstream task model enhancement.
+By generating high-quality data, ProReason effectively distills its superior capabilities into downstream task models, enabling them to demonstrate exceptional performance in complex visual reasoning tasks.
 Reduced information Mission or Redundancy.
-With the close collaboration of the sub-agents in Proactive Visual Perception step, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy.
-Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions.
-This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
+Through sub-agent collaboration in Proactive Visual Perception, ProReason extracts essential visual details via question-oriented manner, preventing information omission or redundancy.
+Meanwhile, the Memory component retains only Vision Expert’s observations and Insight Expert’s conclusions, creating compact representation that minimizes irrelevant information.
+These advantages effectively reduce the overhead when ProReason functions as a reasoning framework.
 4 Experiments
@@
 Datasets.
-To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a)  [^3] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
+To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023)  [^2] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
 To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers.
-For the results using the official MME calculation method, please refer to Table 16, Table 16, Table 18, Table 16 and Table 18.
 , Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a).
-All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Sec.
-A.2.
+All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Section A.2.
 Base Models.
-We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B [^4] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards.
-and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources.
-As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API.
-In contrast, Llama3-LLaVA-NeXT-8B is a fully open-sourced LVLM developed by the LLaVA team based on the Llama-3-8B LLM (Dubey et al., 2024) and CLIP vision encoder (Radford et al., 2021).
-Qwen2.5-72B-Instruct is a robust LLM designed to deliver high-quality instruction following and handle complex tasks.
-It will be utilized for achieving LLM-assisted multi-modal reasoning.
+We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B- Instruct (Bai et al., 2025) [^3] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards.
+to drive ProReason, and Qwen2.5-72B-Instruct (Yang et al., 2024) and Qwen3-32B (Team, 2025) are selected to drive the text-only sub-agents, enabling LLM-assisted visual reasoning.
+In addition Qwen2.5-VL-7B-Instruct is also utilized as a downstream task model, which is trained with the data generated by ProReason.
+These models are chosen for their leading performance and popularity.
 Baselines.
-Besides the most basic method where models are instructed to answer questions directly, we compare ProReason with two categories of peer methods.
-To determine the benefits of proactive information extraction in ProReason, we first consider two SOTA passive visual reasoning methods, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024).
-Additionally, we choose two multi-step reasoning frameworks of LLMs, Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022), to demonstrate the effects of directly migrating LLM solutions to LVLMs.
-  1. Direct. As indicated by the name, models are required to answer questions directly without dedicated prompts. This baseline is set to evaluate the initial performance of base models.
-  2. VDGD. With image caption prefixed to text instruction, VDGD selects tokens that deviate the least from the description using a formula based on Kullback-Leibler divergence.
-  3. CCoT. Given an image and the question, CCOT first generates a scene graph of the image with LVLMs, and then extracts the answer by prompting the LVLMs with the graph.
-  4.
-CoT.
-CoT is an advanced prompting method that encourages LLMs to break complex tasks down into a series of easy steps, which has been applied broadly and verified to boost the reasoning performance remarkably (Chu et al., 2023).
-  5.
-ReAct.
-ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors.
-To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Reasoning Experts, respectively.
-This aligns with our notions for easy understanding, and is shown in Figure 1.
+In addition to the basic method where models are instructed to answer questions directly, we compare ProReason with the following peer methods.
+First, to explore the effectiveness of directly migrating LLM solutions to LVLMs, we select two multi-step reasoning frameworks from LLMs: Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022).
+Second, to evaluate the advantages of proactive information extraction in ProReason, we examine two passive visual reasoning frameworks, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024), which assist LVLM reasoning by extracting image information into text.
+Additionally, we consider R1-Onevision-7B (Yang et al., 2025), an LVLM based on Qwen2.5-VL-7B-Instruct with deep thinking capabilities similar to Deekseek-R1 (Guo et al., 2025) and OpenAI o1444https://openai.com/index/introducing-openai-o1-preview/.
+We compare it with ProReason-Q3, which also exhibits deep thinking abilities.
+Further details on the baselines and implementation are provided in Section A.3.
 Implementation Details.
 The prompt templates for all methods are shown in Figures 9, 10 and 12.
-Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Reasoning Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process.
+Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Insight Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process.
 If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated.
@@
 This setup, refined through multiple trials, is the most effective.
-Additionally, since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini.
-Table 1: Performance of multiple approaches with two base models across four visual reasoning benchmarks.
+Table 1: Performance of multiple approaches with three base models across four visual reasoning benchmarks.
 “Hallu.” is the abbreviation of HallusionBench.
+The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively.
+"Assisted" stands for LLM-assisted reasoning, which involves replacing the textual sub-agents within frameworks with corresponding LLMs, as detailed in Table 14.
 Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
 Model Method Dataset
-MME MMMU MathVista Hallu.
-Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
-VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6)
-CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4)
-CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7)
-ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
-ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1)
-Average 66.30 43.9 36.06 44.16
-GPT-4o-mini Direct 79.2 48.4 53.0 56.0
-VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6)
-CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7)
-CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3)
-ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
-ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9)
-Average 84.63 54.82 52.63 55.4
-Table 2: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.
-The red and blue texts indicate the improvements and reductions relative to the case of using the same method without the assistance of text-only LLMs, respectively.
-Vision Expert Textual Sub-Agents Method Dataset
-MME MMMU MathVista Hallu.
-Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
-ReAct 68.5 46.7 31.7 43.6
-ProReason 71.5 50.5 38.8 50.9
-Qwen2.5-72B- Instruct ReAct 71.0 (+2.5) 50.4 (+3.7) 34.6 (+2.9) 40.4 (-3.2)
-ProReason 81.3 (+9.8) 56.8 (+6.3) 48.8 (+10.0) 52.3 (+1.4)
-GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
-ProReason 84.7 (+13.2) 54.5 (+4.0) 41.7 (+2.9) 53.1 (+2.2)
+MME MMMU MathVista Hallu. Average
+GPT-4o-mini Direct 79.2 48.4 53.0 56.0 59.2
+VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 59.3 (+0.1)
+CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 61.3 (+2.1)
+CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 64.1 (+4.9)
+ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 60.6 (+1.4)
+ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 67.1 (+7.9)
+Llama3-LLaVA-NeXT-8B Direct 61.5 41.8 37.1 45.8 46.6
+VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6) 47.8 (+1.2)
+CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4) 45.9 (-0.7)
+CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7) 44.8 (-1.8)
+ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2) 47.6 (+1.0)
++ 4o-mini Assisted 73.6 (+12.1) 48.4 (+6.6) 36.2 (-0.9) 46.7 (+0.9) 51.2 (+4.6)
++ Qwen72B Assisted 71.0 (+9.5) 50.4 (+8.6) 34.6 (-2.5) 40.4 (-5.4) 49.1 (+2.5)
+ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1) 52.9 (+6.3)
++ 4o-mini Assisted 84.7 (+23.2) 54.5 (+12.7) 41.7 (+4.6) 53.1 (+7.3) 58.5 (+11.9)
++ Qwen72B Assisted 81.3 (+19.8) 56.8 (+15.0) 48.8 (+11.7) 52.3 (+6.5) 59.8 (+13.2)
+Qwen2.5-VL-7B-Instruct Direct 74.2 51.8 63.3 53.8 60.8
+VDGD 74.6 (+0.4) 52.3 (+0.5) 62.1 (-1.2) 53.9 (+0.1) 60.7 (-0.1)
+CCoT 82.7 (+8.5) 52.2 (+0.4) 61.7 (-1.6) 55.8 (+2.0) 63.1 (+2.3)
+CoT 72.3 (-1.9) 53.6 (+1.8) 63.7 (+0.4) 55.9 (+2.1) 61.4 (+0.6)
+ReAct 81.9 (+7.7) 51.6 (-0.2) 60.4 (-2.9) 52.9 (-0.9) 61.7 (+0.9)
++ Qwen72B Assisted 83.1 (+8.9) 52.5 (+0.7) 62.2 (-1.1) 54.5 (+0.7) 63.1 (+2.3)
+ProReason 83.8 (+9.6) 57.0 (+5.2) 63.2 (-0.1) 56.4 (+2.6) 64.6 (+3.8)
++ Qwen72B Assisted 92.7 (+18.5) 64.6 (+12.8) 64.0 (+0.7) 60.6 (+6.8) 70.5 (+9.7)
++ Qwen3 Assisted 90.7 (+16.5) 66.2 (+14.4) 67.2 (+3.9) 59.8 (+6.0) 71.0 (+10.2)
 4.2 Main Results
@@
 As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets.
-In contrast, ProReason consistently outperforms all the other baselines for both Llama3-LLaVA-NeXT-8B and GPT-4o-mini model across all benchmarks, with a peak improvement as 13.2%, demonstrating the superiority and task robustness of ProReason.
-Proactive information acquisition surpasses SOTA passive methods, especially in complex visual reasoning tasks.
+In contrast, ProReason consistently surpasses all other baselines across all benchmarks for every base model, enhancing the average performance of GPT-4o-mini by 7.9%, demonstrating the superiority and task robustness of ProReason.
+Section A.5.7 and Table 13 evaluate ProReason on additional benchmarks, including reasoning and VQA tasks, further proving its strong performance and effectiveness across various multi-modal tasks.
+Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage.
+Table 1 illustrates that CoT, utilizing both capabilities simultaneously, does not exhibit consistent performance enhancements over the "Direct" method when applied to the Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B-Instruct models.
+In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively.
+This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
+Proactive information acquisition surpasses peer passive methods, especially in complex visual reasoning tasks.
 Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities.
@@
 In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions.
-This aligns with out previous obervations in Sec.
-2.2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
-Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage.
-As listed in Table 1, when both capabilities are utilized concurrently, CoT consistently degrades performance compared to the “Direct” method across all benchmarks with Llama3-LLaVA-NeXT-8B model, consistent with the findings in Sec.
-2.1.
-In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively.
-This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, demonstrating the effectiveness of capability decoupling.
-ProReason outperforms ReAct with even less token consumption.
-Given the similarity in multi-step reasoning, we compare ProReason and ReAct in terms of performance and cost.
-Specifically, as an LLM-specific multi-step reasoning framework, ReAct only outperforms the “Direct” method on MME and MMMU, but underperforms on MathVista and HallusionBench, showing inferior performance compared to the consistent improvements of ProReason.
-Furthermore, we compare their average token consumption of GPT-4o-mini.
-As shown in Table 13 of Appendix A, ProReason consumes significantly fewer tokens than ReAct on both MME and MathVista tasks, indicating its higher token efficiency and the importance of a compact Memory in reducing token usage.
-Coupled with better performance, this suggests the superiority of ProReason over LLM-specific ReAct framework.
+This aligns with out previous obervations in Section 2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
 Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance.
-As mentioned in Sec.
-3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs.
-To demonstrate the utility of this advantage, we fix the Vision Expert as Llama3-LLaVA-NeXT-8B, and replace other agents with text-only LLMs.
-As listed in Table 2, with the assistance of either Qwen2.5-72B-Instruct or GPT-4o-mini, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers.
+As mentioned in Section 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs.
+To demonstrate the utility of this advantage, we fix the Vision Expert and replace the textual sub-agents in ProReason with text-only LLMs, according to the configuration in Table 14.
+As listed in Table 1, with the assistance of powerful existing LLMs, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers.
+In addition, by configuring the Summarizer to Qwen3-32B, ProReason acquires the same kind of deep reasoning capability as Deekseek-R1 and OpenAI o1, resulting in a 10.2% average performance improvement for the Qwen2.5-VL-7B-Instruct Vision Expert.
 In contrast, ReAct gains a much smaller improvement.
@@
 Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
-4.3 Relative Importance of Sub-agents
-To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Reasoning Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini.
-Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is.
-Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage.
-The experimental results are presented in Table 13.
-Summarizer is the most crucial sub-agent, closely followed by Referee.
-The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively.
-This highlights the critical function of the Summarizer in integrating all available information to conclude final answers.
-Besides, the substitution of Referee leads to a 10.1% reduction on MMMU.
-Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks.
-The analysis in Section A.3.2 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
-Relatively, Dispatcher and Reasoning Expert are the least essential sub-agents.
-Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents.
-This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Reasoning Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks.
-Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance.
-Even if some error occurs, subsequent iterations can compensate for the missing information.
-In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity.
-Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Reasoning Expert have the least impact.
-4.4 The Critical Implication of Decoupling
-To validate the critical implication of decoupling visual perception and textual reasoning in ProReason while eliminating potential confounding factors from prompt engineering and multiple CoT implementations, we systematically integrate sub-agents through three configurations:
-  1. Merge the Vision Expert and Reasoning Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
-  2. Integrate the Dispatcher, Vision Expert, Reasoning Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
-  3. Fully consolidate all five original sub-agents to demonstrate the pivotal role of decomposing visual reasoning tasks into distinct Proactive Visual Perception and Textual Reasoning phases.
-The merged agents preserve identical prompts and maintain the same task execution procedures as their original counterparts, with their prompts shown in Figures 13, 14, and 15.
-Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
-Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks.
-As shown in Table 13, merging Vision and Reasoning Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs.
-CoT’s 58.5).
-Although agent merging also causes performance declines on MME, the merged versions still outperform CoT.
-Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
+Table 2: Performance of different models across two visual benchmarks.
+Based on the performance of the base model(i.e., Qwen2.5-VL-7B-Instruct), red and blue signify the improvement and degradation, respectively.
+Dataset Model
+Qwen2.5-VL-7B-Instruct R1-Onevision-7B ProReason-VL ProReason-Q3
+MME 72.3 88.1 (+15.8) 91.2 (+18.9) 90.8 (+18.5)
+MMMU(val.) 53.6 53.1 (-0.5) 65.4 (+11.8) 67.4 (+13.8)
+Average 63.0 70.6 (+7.6) 78.3 (+15.3) 79.1 (+16.1)
+Table 3: Average token and time consumption of multiple approaches with GPT-4o-mini model on the MME and MathVista benchmarks.
+Dataset Method GPT-4o-mini
+Direct VDGD CCoT CoT ReAct ProReason
+MME Input 393.9 1020.9 1024.3 403.9 1645.0 1286.8
+Output 5.9 155.6 254.3 103.4 197.0 327.2
+Time(s) 6.1 12.3 12.8 7.3 18.9 18.4
+MathVista Input 368.4 955.5 961.2 375.4 3092.8 2238.6
+Output 51.8 263.3 307.3 479.7 845.1 788.6
+Time(s) 4.2 11.5 12.3 12.0 28.8 24.8
+4.3 Downstream Task Model Enhancement
+High-quality visual reasoning data produced by ProReason significantly improves the performance of downstream task models.
+We select the test set that has no answer in the MMMU dataset and gather reasoning processes paired with corresponding answers on this set, generated by two configurations detailed in Table 14: ProReason + Qwen72B Assisted and ProReason + Qwen3 Assisted.
+After eliminating samples with inconsistent answers between these two configurations, we obtain two filtered datasets each containing 5,980 entries.
+These datasets are then used to fine-tune Qwen2.5-VL-7B-Instruct (Section A.4), yielding two models: ProReason-VL and ProReason-Q3.
+Notably, ProReason-Q3 inherits the deep reasoning capabilities of Qwen3-32B, which incorporates a reasoning process containing a <think>...
+</think> part.
+Table 2 illustrates that while R1-Onevision-7B features deep reasoning capabilities and shows performance enhancements on MME, it fails to demonstrate improvement on the MMMU validation set.
+In contrast, both ProReason-VL and ProReason-Q3 achieve performance gains exceeding 10% on the MMMU validation set and also exhibit significant improvements on MME.
+This indicates that ProReason effectively transfers the exceptional reasoning capabilities of LLM-integrated decoupled systems to downstream task models, highlighting the potential for leveraging the robust abilities of existing LLMs to continuously enhance the performance of LVLMs.
+4.4 Efficiency and Complexity Analysis
+ProReason improves both efficiency and performance compared to the baseline method.
+Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets.
+Notably, ProReason requires significantly fewer tokens than ReAct, while achieving superior performance over ReAct as analyzed in Section 4.2.
+Moreover, compared to the visual reasoning frameworks VDGD and CCoT, which also involve multiple image inputs, ProReason’s token consumption on MME is only about 20% higher, yet it achieves a 12.7% performance improvement.
+Regarding time efficiency, ProReason achieves 11.1% higher performance than VDGD and CCoT on MME with comparable time consumption (18.4s/sample vs.
+VDGD’s 12.3s and CCoT’s 12.8s), while ReAct is slower than ProReason across datasets, highlighting ProReason’s efficiency advantage.
+Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
+ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms.
+The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency.
+Additionally, ProReason dynamically adjusts Proactive Visual Perception iterations based on question difficulty, minimizing overhead for simple tasks while enhancing complex problem-solving.
+As shown in Table 12, MathVista’s greater challenge prompts more iterations compared to MME, resulting in higher token usage and longer reasoning times.
+This aligns with the difficulty levels of the datasets, demonstrating the adaptive nature of ProReason.
+4.5 Ablation Study and Further Analysis
+Relative Importance of Sub-agents.
+To evaluate the importance of each sub-agent in ProReason, in Section A.5.1, we replace Dispatcher, Vision Expert, Insight Expert, Referee, and Summarizer individually with the less capable Llama3-LLaVA-NeXT-8B(which demonstrates weaker visual understanding and textual reasoning capabilities), while keeping the other sub-agents as GPT-4o-mini.
+The performance degradation on the MME and MMMU benchmarks is then used to measure the significance of each sub-agent.
+Results indicate that the Summarizer is the most critical sub-agent, closely followed by Referee.
+Which One is More Crucial: Visual Understanding or Textual Reasoning
+In Section A.5.2, we perform comparative experiments by substituting the vision expert and text sub-agents in the ProReason with Llama3-LLaVA-NeXT-8B and GPT-4o-mini respectively.
+Our findings indicate that while both visual understanding and textual reasoning capabilities are essential for multimodal tasks, textual reasoning ability holds greater significance in visual reasoning tasks.
+This result is consistent with our earlier analysis in Section A.5.1, which identifies the Summarizer and Referee as the most critical sub-agents.
+The Critical Implication of Decoupling.
+In Section A.5.3, to validate the necessity of decoupling visual perception and textual reasoning in ProReason, we systematically merge sub-agents through three configurations while preserving identical prompts and procedures, with observed performance degradation quantitatively demonstrating the critical role of decomposed processing in enhancing capabilities.
+The experimental results indicate that Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks
+Reasoning Process Evaluation of ProReason.
+In Section A.5.4, we evaluate the responses generated by ProReason using LLMs.
+The analysis reveals that, compared to CoT, ProReason produces more relevant answers with reduced redundancy and deficiency, consistent with its enhanced performance.
+Referee’s Dispel of Hallucinations.
+In Section A.5.5, adhering to the settings in Section 4.1, we assess ProReason (powered by GPT-4o-mini) on MMMU and HallusionBench with different attempt allowances (1/3/5).
+When attempts are unsuccessful, systematic Memory clearance is triggered (determined by the Referee’s lack of sufficient information).
+As the number of attempts increases, the Referee has more opportunities for information filtering.
+The observed performance variations illustrate the crucial influence of the Referee’s decision - making and filtering efficiency on the system’s capabilities.
+Experimental results indicate that the Referee module effectively filters hallucinated information to improve the visual comprehension capabilities of our framework.
+Frequency of selection of various experts.
+In Section A.5.6, we assess how often the Dispatcher selects the Vision Expert or Insight Expert across both MME and MMMU benchmarks.
+The experimental results demonstrate that ProReason adaptively adjusts the frequencies of expert selection, leading to consistent performance improvements.
 5 Conclusion
@@
 Contradictory Information among Agents.
-Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and reasoning expert.
-However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the reasoning expert) tend to adhere to the available information instead of questioning it.
+Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and Insight Expert.
+However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the Insight Expert) tend to adhere to the available information instead of questioning it.
 This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
+We also examined existing multi-step reasoning frameworks,
+such as ReAct (Yao et al., 2022), ToT (Yao et al., 2024), and Insight-V (Dong et al., 2024), aiming to find solutions for resolving accumulated errors and contradictory information.
+However, we found that these approaches may also fail to identify effective solutions and do not discuss the impact of the aforementioned drawbacks.
+We sincerely WELCOME any constructive discussions regarding accumulated errors and contradictory information!
+Besides, ProReason demonstrates significantly improved overall performance, suggesting fewer errors made by ProReason in the whole task distribution.
 7 Ethics Statement
@@
 Large Visual-Language Model.
-Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents.
-In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024b) families, have been extensively developed.
-In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success.
+Recently, large vision-language models (LVLMs) (Yang et al., 2024; Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents.
+In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024c) families, have been extensively developed.
+In the closed-source domain, proprietary models such as GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ and Gemini Pro 2.5666https://deepmind.google/technologies/gemini/pro/ have also achieved significantly success.
+Additionally, multi-agent frameworks like VipAct (Zhang et al., 2024c) have been developed to improve LVLMs’ perception of visual details.
 Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously.
-This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Ghosh et al., 2024) and motivating more effective solutions.
-Passive Visual Reasoning.
-Extracting information from images into text can effectively assist LVLM in performing visual reasoning tasks.
-Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) first describes the image before prefixing this description to the prompt, assisting LVLMs on visual reasoning tasks.
-Furthermore, Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) directs LVLMs to generate scene graphs (SGs) that serve as a bridge between the visual and textual domains, aiding LVLMs in subsequent tasks.
-However, most of these methods employ a question-agnostic and reasoning-free visual extraction process, where image descriptions are not tailored to specific questions, and no reasoning is applied to infer additional information for improved descriptions.
-These “passive” approaches lead to the inclusion of irrelevant or redundant information, ultimately degrading performance.
-In contrast, ProReason adopts question-oriented agents to collect necessary and sufficient information, effectively circumventing these drawbacks.
+This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Liu et al., 2023a; Ghosh et al., 2024) and motivating more effective solutions.
 Multi-step Reasoning Framework.
-Multi-step reasoning frameworks have been developed for LLMs to achieve better performance by breaking down complex questions into easier ones (Pan et al., 2024).
-As a representative method, Chain-of-Thought (CoT) (Wei et al., 2022) enhances the arithmetic and commonsense reasoning capabilities by explicitly generating intermediate reasoning steps before concluding the final answers.
-Tree-of-Thoughts (ToT) (Yao et al., 2024) further refines the CoT mechanism by allowing LLMs to consider multiple reasoning paths and do self-assessment before making decisions.
-Considering that the inherent knowledge of LLMs may not be sufficient to complete tasks, ReAct (Yao et al., 2022) integrates information retrieval into the reasoning chains, enabling models to pause to verify results and determine whether additional information is needed before proceeding.
-Nevertheless, multi-step reasoning frameworks designed for text-only LLMs are not fully applicable to visual reasoning tasks, and may even impair the performance of LVLMs (Ghosh et al., 2024).
+Multi-step reasoning frameworks improve LLM performance by breaking down complex tasks.
+Chain-of-Thought (CoT) (Wei et al., 2022) enhances reasoning via explicit intermediate steps, demonstrating effectiveness in both textual and visual tasks (Zhang et al., 2024b), while Tree-of-Thoughts (ToT) (Yao et al., 2024) extends this by evaluating multiple reasoning paths.
+ReAct (Yao et al., 2022) integrates dynamic knowledge retrieval during reasoning.
+In visual reasoning, several frameworks assist LVLM by extracting image information into text.
+Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) describes the image and appends this description to the prompt, aiding LVLMs in visual reasoning tasks.
+Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) guides LVLMs to create scene graphs (SGs) that link visual and textual domains, supporting subsequent tasks.
+However, these methods often use a question-agnostic, reasoning-free visual extraction process, resulting in irrelevant or redundant information.
+Insight-V (Dong et al., 2024) trains LVLM with a multi-agent system, but all agents rely on a single LVLM, blending visual understanding with textual reasoning and failing to address the challenge of effectively managing both.
+Unfortunately, this method does not release its code and prompts, resulting in low reproducibility and making comparison difficult.
+In response to these drawbacks, we introduce ProReason, which decouples visual reasoning tasks into proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and makes all prompts available.
+By leveraging the strengths of the decoupled system, ProReason effectively integrates existing powerful LLMs to achieve high-performance visual reasoning and successfully transfers this capability to downstream task models.
 A.2 Dataset
-To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a).
+To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a).
 Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
@@
 It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
-A.3 Supplementary Results and Analysis
-A.3.1 Which One is More Crucial: Visual Understanding or Textual Reasoning?
+A.3 Baselines
+  1. Direct. As indicated by the name, models are required to answer questions directly without dedicated prompts. This baseline is set to evaluate the initial performance of base models.
+  2.
+CoT.
+CoT is an advanced prompting method that encourages LLMs to break complex tasks down into a series of easy steps, which has been applied broadly and verified to boost the reasoning performance remarkably (Chu et al., 2023).
+  3.
+ReAct.
+ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors.
+To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Insight Experts, respectively.
+This aligns with our notions for easy understanding, and is shown in Figure 1.
+  4.
+VDGD.
+VDGD involves two main steps: initially, LVLMs generate detailed image captions, which are then incorporated into prompts to aid inference.
+During the inference process, VDGD also utilizes a formula based on Kullback-Leibler divergence to select tokens that minimally deviate from the description, thereby enhancing the relevance of the model’s reasoning to the image [^7] Since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini..
+  5. CCoT. Given an image and the question, CCOT first generates a scene graph of the image with LVLMs, and then extracts the answer by prompting the LVLMs with the graph.
+  6.
+R1-Onevision.
+R1-Onevision-7B is an advanced multimodal reasoning model based on Qwen2.5-VL-7B-Instruct that has deep thinking capabilities akin to Deepseek-R1 by transforming images into structured textual representations and employing a training framework that merges supervised fine-tuning with reinforcement learning.
+A.4 Model Training
+[…]
+A.5 Supplementary Results and Analysis
+Table 4: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
+Model Method
+Direct CoT VDGD
+Llama3-LLaVA-NeXT-8B 41.8 41.5 42.7
+Qwen2.5-VL-7B-Instruct 51.8 52.2 52.6
+Table 5: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion.
+“True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
+Score Llama3-LLaVA-NeXT-8B
+True False
+Detail Level 4.43 3.93
+Question Relevance 3.87 3.30
+Reasoning Effective Info Inclusion 3.91 3.57
+A.5.1 Relative Importance of Sub-agents
+To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Insight Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini.
+Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is.
+Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage.
+The experimental results are presented in Table 6.
+Summarizer is the most crucial sub-agent, closely followed by Referee.
+The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively.
+This highlights the critical function of the Summarizer in integrating all available information to conclude final answers.
+Besides, the substitution of Referee leads to a 10.1% reduction on MMMU.
+Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks.
+The analysis in Section A.5.5 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
+Relatively, Dispatcher and Insight Expert are the least essential sub-agents.
+Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents.
+This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Insight Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks.
+Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance.
+Even if some error occurs, subsequent iterations can compensate for the missing information.
+In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity.
+Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Insight Expert have the least impact.
+Table 6: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
+For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
+The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
+Model Agent Dataset
+MME MMMU
+GPT-4o-mini 90.4 61.6
+Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
+Vision Expert 84.7 (-5.7) 54.5 (-7.1)
+Insight Expert 88.7 (-1.7) 60.2 (-1.4)
+Referee 89.6 (-1.1) 51.5 (-10.1)
+Summarizer 84.2 (-6.2) 51.0 (-10.6)
+A.5.2 Which One is More Crucial: Visual Understanding or Textual Reasoning?
 ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs.
@@
 However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert.
-This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec.
-4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
+This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Section A.5.1 that identifies the Summarizer and Referee as the most crucial sub-agents.
 Table 7: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks.
@@
 Direct 61.5 41.8
-A.3.2 Referee’s Dispel of Hallucinations
+A.5.3 The Critical Implication of Decoupling
+To validate the critical implication of decoupling visual perception and textual reasoning in ProReason while eliminating potential confounding factors from prompt engineering and multiple CoT implementations, we systematically integrate sub-agents through three configurations:
+  1. Merge the Vision Expert and Insight Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
+  2. Integrate the Dispatcher, Vision Expert, Insight Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
+  3. Fully consolidate all five original sub-agents to demonstrate the pivotal role of decomposing visual reasoning tasks into distinct Proactive Visual Perception and Textual Reasoning phases.
+The merged agents preserve identical prompts and maintain the same task execution procedures as their original counterparts, with their prompts shown in Figures 13, 14, and 15.
+Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
+Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks.
+As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs.
+CoT’s 58.5).
+Although agent merging also causes performance declines on MME, the merged versions still outperform CoT.
+Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
+Table 8: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments.
+The blue text highlights the performance degradation due to sub-agent integration.
+GPT-4o-mini Dataset
+MMMU MME
+ProReason 61.6 91.9
+CoT 58.5 87.8
+Vision & Reasoning
+Expert Integration
+58.6 (-3.0) 90.4 (-1.5)
+Dispatcher,
+Vision & Insight Expert,
+& Referee Integration
+57.8 (-3.8) 90.8 (-1.1)
+All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
+A.5.4 Reasoning Process Evaluation of ProReason
+Inspired by Liu et al.
+(2023c), we design a pipeline using LLMs to further analyze the performance of ProReason.
+Given that GPT-4o achieved a 70.3% score on MMMU, significantly surpassing GPT-4o-mini’s 58.5% and ProReason driven by GPT-4o-mini’s 61.6%, we adopt GPT-4o’s CoT-based answers on MMMU as the standard answer.
+Using this standard answer, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason powered by GPT-4o-mini (ProReason-Memory), and the reasoning process of ProReason’s Summarizer (ProReason-Summarizer) on MMMU.
+This assessment focuses on three key metrics: the relevance to standard answer (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement.
+The evaluation process is driven by GPT-4o and the prompt is shown in Figure 12.
+Specifically, compared to GPT-4o-mini (CoT), ProReason-Summarizer produces more relevant answers with less redundancy and deficiency, aligning with its improved performance.
+Compared to ProReason-Summarizer, ProReason-Memory exhibits the same RE, higher RI and lower MI scores.
+This suggests that ProReason allows some redundancy to prevent information loss in memory, as the former typically leads to more serious consequences than the latter.
+Subsequently, Summarizer can leverage its powerful reasoning capabilities to select the most relevant memory.
+Table 9: Performance of ProReason driven by GPT-4o-mini assessed by LLMs compared to CoT on MMMU benchmark.
+Mainly includes three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI ↓), where arrows indicate the directions of improvement.
+GPT-4o-mini Metrics
+RE ↑ RI ↓ MI ↓
+COT 4.67 3.33 1.40
+ProReason-Memory 4.83 3.66 1.17
+ProReason-Summarizer 4.83 2.88 1.33
+A.5.5 Referee’s Dispel of Hallucinations
 Following the implementation details outlined in Section 4.1, we evaluate the ProReason system powered by GPT-4o-mini on both MMMU and HallusionBench under varying attempt allowances (i.e., 1, 3, and 5 attempts).
@@
 HallusionBench 57.1 58.9 59.9
-A.3.3 Performance of ProReason on VQA task
-ProReason also excels in conventional VQA tasks.
-We perform an evaluation on the VQA benchmark A-OKVQA dataset (Schwenk et al., 2022).
-The visual complexity of A-OKVQA is comparable to MME, but its reasoning requirements are less demanding.
-Table 5 illustrates the performance of ProReason, powered by GPT-4o-mini, on the A-OKVQA, MME, and MMMU benchmarks.
-ProReason achieves the highest performance across all three benchmarks, indicating the positive impact of ProReason on any form of reasoning-involved multi-modal tasks.
-ProReason is particularly effective on datasets with greater complexity due to its decoupled capabilities and LLM-assisted reasoning.
-Table 5 demonstrates that ProReason outperforms the "Direct" method by 2.7%, 11.2%, and 13.2% on A-OKVQA, MME, and MMMU, respectively, as the visual and reasoning complexity increases.
-This underscores the advantage of ProReason in handling more challenging tasks, making it a robust choice for complex problem-solving scenarios.
-Table 5: Performance of multiple approaches with GPT-4o-mini across three visual benchmarks.
-GPT-4o-mini Dataset
-A-OKVQA MME MMMU
-Direct 78.9 79.2 48.4
-VDGD 79.4 82.3 51.4
-CCoT 79.2 80.8 54.2
-COT 80.9 87.8 58.5
-ReAct 80.6 87.3 54.8
-ProReason 81.3 (+2.7) 90.4 (+11.2) 61.6 (+13.2)
-A.3.4 Evaluation of ProReason based on LLMs
-In Sec.
-2.1, we utilize a pipeline based on LLMs to evaluate the correlation between model responses and image content.
-We aim to apply this effective method once more to assess the performance of ProReason.
-Initially, we adopted the reasoning chains generated by GPT-4o-mini (CoT) as the standard answers, given its superior performance with a score of 58.5 on the MMMU benchmark, compared to other models.
-However, due to ProReason’s enhancement of GPT-4o-mini’s score to 61.6, the use of GPT-4o-mini’s answers became inadequate.
-Consequently, we opted for GPT-4o to generate the standard answers, achieving a score of 69.1 on MMMU.
-As shown in Table 6, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason (ProReason-Memory), and the reasoning process of Summarizer (ProReason-Summarizer), which generates the final answer of ProReason with the prompt template in Figure 12.
-This assessment focuses on three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement.
-Specifically, compared to GPT-4o-mini (CoT), ProReason-Summarizer produces more relevant answers with less redundancy and deficiency, aligning with its improved performance.
-Compared to ProReason-Summarizer, ProReason-Memory exhibits the same RE, higher RI and lower MI scores.
-This suggests that ProReason allows some redundancy to prevent information loss in memory, as the former typically leads to more serious consequences than the latter.
-Subsequently, Summarizer can leverage its powerful reasoning capabilities to select the most relevant memory.
-Table 6: Performance of ProReason driven by GPT-4o-mini assessed by LLMs compared to CoT on MMMU benchmark.
-Mainly includes three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI ↓), where arrows indicate the directions of improvement.
-GPT-4o-mini Metrics
-RE ↑ RI ↓ MI ↓
-COT 4.67 3.33 1.40
-ProReason-Memory 4.83 3.66 1.17
-ProReason-Summarizer 4.83 2.88 1.33
-A.3.5 Frequency of selection of various experts
-As listed in Table 7, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities.
-Specifically, compared to MME, the frequencies for both the Vision and Reasoning Experts are higher on the MMMU benchmark, aligning with their difficulty levels.
+A.5.6 Frequency of selection of various experts
+As listed in Table 11, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities.
+Specifically, compared to MME, the frequencies for both the Vision and Insight Experts are higher on the MMMU benchmark, aligning with their difficulty levels.
 Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%).
-Despite the lower frequency of the reasoning expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions.
+Despite the lower frequency of the Insight Expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions.
 Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
-Table 7: Frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks.
+Table 11: Frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks.
 Dataset GPT-4o-mini
-Vision Expert Reasoning Expert
+Vision Expert Insight Expert
 MME 1.16 0.12
 MMMU 1.64 0.38
-Table 8: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
-Model Method
-Direct CoT VDGD
-Llama3-LLaVA-NeXT-8B 41.8 41.5 42.3
-LLaVA-OneVision-Qwen2-7B-OV 49.1 46.5 49.7
-Qwen-VL-Chat 36.2 35.4 37.1
-Table 9: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
-Model Image Relevance Score
-True False
-Llama3-LLaVA-NeXT-8B 3.70 3.36
-LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
-Qwen-VL-Chat 3.63 2.80
-Table 10: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion.
-“True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
-Score Llama3-LLaVA-NeXT-8B
-True False
-Detail Level 4.43 3.93
-Question Relevance 3.87 3.30
-Reasoning Effective Info Inclusion 3.91 3.57
-Table 11: Average token consumption of ProReason and ReAct with GPT-4o-mini model on the MME and MathVista benchmarks.
-Method MME MathVista
-Input Output Input Output
-ProReason 1286.8 327.2 2238.6 788.6
-ReACT 1645.0 197.0 3092.8 845.1
-Table 12: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
-For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
-The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
-Model Agent Dataset
-MME MMMU
-GPT-4o-mini 90.4 61.6
-Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
-Vision Expert 84.7 (-5.7) 54.5 (-7.1)
-Reasoning Expert 88.7 (-1.7) 60.2 (-1.4)
-Referee 89.6 (-1.1) 51.5 (-10.1)
-Summarizer 84.2 (-6.2) 51.0 (-10.6)
-Table 13: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments.
-The blue text highlights the performance degradation due to sub-agent integration.
-GPT-4o-mini Dataset
-MMMU MME
-ProReason 61.6 91.9
-CoT 58.5 87.8
-Vision & Reasoning
-Expert Integration
-58.6 (-3.0) 90.4 (-1.5)
-Dispatcher,
-Vision & Reasoning Expert,
-& Referee Integration
-57.8 (-3.8) 90.8 (-1.1)
-All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
-Table 14: Performance of multiple approaches with two base models across four visual reasoning benchmarks. Different from Table 1, MME scores are calculated with the official method here.
+A.5.7 Performance Evaluation of ProReason on Supplementary Datasets
+To further assess the effectiveness of ProReason across diverse visual tasks, we expand upon the four datasets introduced in Section 4.1 by incorporating three additional benchmarks: MathVerse (Zhang et al., 2024a), MMStar (Chen et al., 2024a), and A-OKVQA (Schwenk et al., 2022).
+MathVerse focuses on visual reasoning within mathematical contexts, while MMStar highlights visual dependency and data reliability.
+A-OKVQA, in contrast, serves as a knowledge-based VQA dataset requiring minimal reasoning.
+For MathVerse, we utilize the Text Lite subset, which prioritizes visual reasoning by minimizing textual content.
+In MMStar, subsets related to coarse and fine-grained perception that are unrelated to reasoning are excluded.
+Table 13 presents ProReason’s performance on each dataset, showcasing significant improvements across various forms of multi-modal tasks and highlighting the effectiveness and generalizability of the perception and reasoning decoupling approach.
+Table 12: Number of iterations required for each problem during the Proactive Visual Perception phase on MME and MathVista benchmarks.
+Model Dataset
+MME MathVista
+GPT-4o-mini 1.28 2.13
+Table 13: Performance of multiple approaches with GPT-4o-mini across 7 visual benchmarks.
+“Hallu.” is the abbreviation of HallusionBench.
+Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
 Model Method Dataset
-MME MMMU MathVista HallusionBench
-Llama3-LLaVA- NeXT-8B Direct 35.9 41.8 37.1 45.8
-VDGD 37.4 42.3 36.1 44.2
-CCoT 35.4 40.5 36.8 37.4
-CoT 25.8 41.5 35.9 43.1
-ReAct 38.5 46.7 31.7 43.6
-ProReason 41.6 50.5 38.8 50.9
-GPT-4o-mini Direct 55.2 48.4 53.0 56.0
-VDGD 58.4 51.4 51.2 52.4
-CCoT 53.5 54.2 53.6 56.7
-CoT 72.9 58.5 53.8 56.3
-ReAct 71.2 54.8 49.3 51.1
-ProReason 77.9 61.6 54.9 59.9
-Table 15: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks.
-Different from Table 2, MME scores are calculated with the official method here.
-Vision Expert Textual Sub-Agents Method Dataset
-MME MMMU MathVista Hallu.
-Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B ReAct 385 46.7 31.7 43.6
-ProReason 416 50.5 38.8 50.9
-Qwen2.5-72B- Instruct ReAct 428 50.4 34.6 40.4
-ProReason 603 56.8 48.8 52.3
-GPT-4o-mini ReAct 473 48.4 36.2 46.7
-ProReason 613 54.5 41.7 53.1
-Table 16: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks.
-Different from Table 3, MME scores are calculated with the official method here.
-Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
-Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
-MME 613 512 416 286 359
-MMMU 54.5 53.4 52.5 41.5 41.8
-Table 17: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks.
-For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini.
-Different from Table 13, MME scores are calculated with the official method here.
-Model Agent Dataset
-MME MMMU
-GPT-4o-mini 779 61.6
-Llama3- LLaVA- NeXT-8B Dispatcher 746 60.9
-Vision Expert 613 54.5
-Reasoning Expert 743 60.2
-Referee 762 51.5
-Summarizer 641 51.0
-Table 18: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments.
-Different from Table 13, MME scores are calculated with the official method here.
-GPT-4o-mini Dataset
-MMMU MME
-ProReason 61.6 779
-CoT 58.5 729
-Vision & Reasoning
-Expert Integration
-58.6 673
-Dispatcher,
-Vision & Reasoning Expert,
-& Referee Integration
-57.8 714
-All Sub-Agents Integration 56.8 711
-A.4 Demonstrative Examples
+MME MMMU MathVista Hallu. MathVerse MMStar A-OKVQA
+GPT-4o-mini Direct 79.2 48.4 53.0 56.0 28.2 46.2 78.6
+VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 30.1 (+1.9) 47.0 (+0.8) 79.4 (+0.8)
+CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 29.2 (+1.0) 45.4 (-0.8) 79.2 (+0.6)
+CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 28.9 (+0.7) 47.2 (+1.0) 80.9 (+2.3)
+ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 30.4 (+2.2) 46.7 (+0.5) 80.6 (+2.0)
+ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 31.6 (+3.4) 49.1 (+2.9) 81.3 (+2.7)
+Average 84.63 54.82 52.63 55.4 29.7 46.9 80.0
+Table 14: Different model configurations of Textual Sub-Agents for implementing LLM-assisted visual reasoning in ProReason.
+The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively.
+"Assisted" stands for LLM-assisted reasoning.
+Model Method Vision Expert Summarizer Other Textual Sub-Agents
+Llama3-LLaVA-NeXT-8B ProReason + 4o-mini Assisted Llama3-LLaVA-NeXT-8B GPT-4o-mini GPT-4o-mini
+ProReason + Qwen72B Assisted Llama3-LLaVA-NeXT-8B Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
+Qwen2.5-VL-7B-Instruct ProReason + Qwen72B Assisted Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
+ProReason + Qwen3 Assisted Qwen2.5-72B-Instruct Qwen3-32B Qwen2.5-72B-Instruct
+A.6 Demonstrative Examples
 Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason.
@@
 The vision expert incorrectly provided the coordinates of the three vertices of the triangle, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
-A.5 Prompts
+A.7 Prompts
 Figure 8: Prompt templates of Relevance Score and caption effectiveness evaluation.
@@
 Figure 10: Prompt templates of Dispatcher.
-Figure 11: Prompt templates of Vision Expert, Reasoning Expert, Referee, and Summarizer.
+Figure 11: Prompt templates of Vision Expert, Insight Expert, Referee, and Summarizer.
 Figure 12: Prompt templates of LLMs-based ProReason evaluation.
-Figure 13: Prompt templates of Vision and Reasoning Expert Integration.
-Figure 14: Prompt templates of Dispatcher, Vision Expert Reasoning Expert and Referee Integration.
+Figure 13: Prompt templates of Vision and Insight Expert Integration.
+Figure 14: Prompt templates of Dispatcher, Vision Expert Insight Expert and Referee Integration.
 Figure 15: Prompt templates of All Five Sub-Agent Integration.

### v5

@@
 Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
+The code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
 ProReason: Multi-Modal Proactive Reasoning
 with Decoupled Eyesight and Wisdom
-Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Lui, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, u3638070@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
+Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Liu, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, {u3011211, u3009618, u3013005, u3638070}@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
 1 Introduction
 In recent years, large language models (LLMs) (Yang et al., 2024; Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024).
-This progress has also sparked interest in developing large vision-language models (LVLMs) (Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks.
+This progress has also sparked interest in developing large vision-language models (LVLMs) (Xiaomi and Team, 2025; Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks.
 However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences.
@@
 MME MMMU MathVista Hallu. Average
-GPT-4o-mini Direct 79.2 48.4 53.0 56.0 59.2
+GPT-4o-mini
+Direct 79.2 48.4 53.0 56.0 59.2
 VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 59.3 (+0.1)
@@
 This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
-Proactive information acquisition surpasses peer passive methods, especially in complex visual reasoning tasks.
+Proactive information acquisition surpasses peer passive methods
+, especially in complex visual reasoning tasks.
 Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities.
@@
 4.4 Efficiency and Complexity Analysis
-ProReason improves both efficiency and performance compared to the baseline method.
+ProReason improves both efficiency and performance compared to the baseline method
+.
 Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets.
@@
 Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
-ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms.
+ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms
+.
 The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency.
@@
 Furthermore, we will release the code upon publication to facilitate easy public reproducibility.
+8 Acknowledgements
+We want to thank our anonymous AC and reviewers for their feedback.
+This work was supported in part by Hong Kong Innovation and Technology Commission’s Innovation and Technology Fund (Award No.
+ITS/269/22FP), Hong Kong RGC grants C7004-22G (CRF) and CRS_PolyU501/23 (CRS).
 Appendix A Appendix
@@
 A.4 Model Training
-[…]
+we use Supervised Fine-Tuning(SFT) and employ the parameter-efficient fine-tuning method LoRA (Hu et al., 2021).
+Specifically, we uniformly set the learning rate to $1\times 10^{-4}$, $\texttt{lora\_dropout}=0$, and train the ProReason-VL for 1 epoch and ProReason-Q3 for 2 epoch.
+These parameters are the optimal values obtained after multiple attempts.
 A.5 Supplementary Results and Analysis
@@
 GPT-4o-mini 90.4 61.6
-Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
+Llama3-
+LLaVA-
+NeXT-8B
+Dispatcher 88.8 (-1.6) 60.9 (-0.7)
 Vision Expert 84.7 (-5.7) 54.5 (-7.1)
@@
 MME MMMU
-GPT-4o -mini Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
+GPT-4o
+-mini
+Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
 Vision Expert 77.8 (+6.3) 53.4 (+0.9)
-Llama3- LLaVA- NeXT-8B All Sub-Agents 71.5 52.5
+Llama3-
+LLaVA-
+NeXT-8B
+All Sub-Agents 71.5 52.5
 COT 58.8 41.5
@@
 Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
-Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks.
+Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks
+.
 As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs.
@@
 The observed performance variations of ProReason across different attempt quotas demonstrate the critical impact of Referee’s decision-making mechanism and information filtering efficacy on system capability.
-The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework.
+The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework
+.
 As demonstrated in Table 10, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5.
@@
 Table 13 presents ProReason’s performance on each dataset, showcasing significant improvements across various forms of multi-modal tasks and highlighting the effectiveness and generalizability of the perception and reasoning decoupling approach.
+A.5.8 The Impact of Iterative Reasoning on ProReason’s Performance
+As shown in the Table 14, we evaluated the performance of ProReason, powered by GPT-4o-mini, on the MME, MMMU, and MathVista datasets, given different maximum loop iterations.
+The performance of ProReason is observed to enhance, as the maximum iterations increase up to three, after which the performance tends to stabilize.
+This observation indicates that most questions in the MME, MMMU, and MathVista benchmarks can be effectively addressed within three cycles of question-oriented visual information extraction.
+In cases of a few difficult questions, ProReason can adaptively extend the number of iterations, proactively gathering more information from the image to arrive at the solution.
 Table 12: Number of iterations required for each problem during the Proactive Visual Perception phase on MME and MathVista benchmarks.
@@
 Average 84.63 54.82 52.63 55.4 29.7 46.9 80.0
+Table 14: Performance of ProReason with GPT-4o-mini across MME, MMMU, and MathVista benchmarks, given different maximum loop iterations.
+The numbers in parentheses signify the improvement or degradation over the direct method, with red for improvement and blue for degradation.
+Dataset Direct VDGD CCoT CoT ReAct ProReason (max loop iterations)
+1 2 3 4 5
+MME 79.2 82.3 (+3.1) 80.8 (+1.6) 87.8 (+8.6) 87.3 (+8.1) 90.8 (+11.6) 91.5 (+12.3) 91.2 (+12.0) 91.5 (+12.3) 91.9 (+12.7)
+MMMU 48.4 51.4 (+3.0) 54.2 (+5.8) 58.5 (+10.1) 54.8 (+6.4) 59.2 (+10.8) 60.8 (+12.4) 61.3 (+12.9) 61.1 (+12.7) 61.6 (+13.2)
+MathVista 53.0 51.2 (-1.8) 53.6 (+0.6) 53.8 (+0.8) 49.3 (-3.7) 52.3 (-0.7) 53.8 (+0.8) 54.2 (+1.2) 54.5 (+1.5) 54.9 (+1.9)
 Table 15: Different model configurations of Textual Sub-Agents for implementing LLM-assisted visual reasoning in ProReason.
@@
 "Assisted" stands for LLM-assisted reasoning.
-Model Method Vision Expert Summarizer Other Textual Sub-Agents
-Llama3-LLaVA-NeXT-8B ProReason + 4o-mini Assisted Llama3-LLaVA-NeXT-8B GPT-4o-mini GPT-4o-mini
+Model
+Method Vision Expert Summarizer Other Textual Sub-Agents
+Llama3-LLaVA-NeXT-8B
+ProReason + 4o-mini Assisted Llama3-LLaVA-NeXT-8B GPT-4o-mini GPT-4o-mini
 ProReason + Qwen72B Assisted Llama3-LLaVA-NeXT-8B Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct